### v2

# On stabilizers in finite permutation groups

Abstract.
Let $G$ be a permutation group on the finite set $\Omega$. We prove various results about partitions of $\Omega$ whose stabilizers have good properties. In particular, we obtain the following three theorems: In every solvable permutation group, there is a subset whose stabilizer has all orbits of length at most $6$, which is best possible and answers two questions of Babai. Every solvable maximal subgroup of any almost simple group has derived length at most $10$, which is best possible. In every primitive group with solvable stabilizer, there are two points whose pointwise stabilizer has derived length bounded by an absolute constant.
Key words and phrases:
Permutation groups, primitive groups, asymmetry, derived length
2020 Mathematics Subject Classification:
Primary: 20B05, 20D05, 05A18
1. Introduction
Let $G$ be a permutation group on the finite set $\Omega$. One of the most studied topics in permutation group theory is finding a partition of $\Omega$ whose stabilizer is trivial. For example, a subset $\Delta\subseteq\Omega$ is said to be in a regular orbit if its setwise stabilizer is trivial. In general, the existence of such a $\Delta$ is far from certain [Glu83, Ser97], and it is necessary to stabilize many subsets. On the other hand, the base size is the minimum number of points such that their pointwise stabilizer is trivial. This corresponds to fixing a partition made by some single points and the complement. Finding upper bounds for the base size is a difficult problem that has produced an immense volume of research, and some of the most impressive results obtained are [Bab81, Pyb93, Ser96, DHM18]. For more details on these topics, we refer the reader to the pleasant survey [BC11].
In this paper, we are interested in a small variation on these questions. More precisely, we fix only the bare minimum (so one subset, or a pair of points) and we do not aim to find a trivial stabilizer, since this would be impossible in general, but to find a stabilizer that has at least some good properties. We will see that this can be done in a few interesting situations. Our first main result is the following:
Theorem 1.1.
Let $G\leqslant\mathop{\mathrm{Sym}}(\Omega)$ be a solvable permutation group. Then there exists a subset of $\Omega$ whose setwise stabilizer has all orbits of length at most $6$.
The following consequence is immediate.
Corollary 1.2.
Let $G\leqslant\mathop{\mathrm{Sym}}(\Omega)$ be a solvable permutation group. Then there exists a subset of $\Omega$ whose setwise stabilizer has derived length at most $3$.
For primitive solvable groups, we can find a subset whose stabilizer has orbits of length at most $3$ and is metabelian. In the beginning we hoped to prove the same for all transitive solvable groups, but this happened to be false (see Examples 2.8 and 2.9). In fact, the results above are the best possible, and answer two questions of Babai [Bab22, Questions 7a-7b]. Theorem 1.1 seems to be the first statement concerning stabilizers of imprimitive groups where the property involved is not preserved by group extensions, and part of the difficulty of the proof is in finding the right statement to prove by induction. [^1] Babai [Bab22, Theorem 12.1] proves the same result with a nonexplicit constant.
We come to the question of fixing two points. Extending the work of Seress [Ser96], Burness [Bur21] showed that in a primitive group with solvable stabilizer, there always exist $5$ points whose pointwise stabilizer is trivial. For primitive solvable groups, in [LS24] the authors observed that there exist two points whose pointwise stabilizer has bounded derived length. On the other hand, motivated by various applications, Li and Zhang [LZ11] have classified the solvable maximal subgroups of the almost simple groups (see also [Bur21]). Combining their work with some properties of the finite simple groups, we remark the following:
Theorem 1.3.
Let $G$ be an almost simple group, and let $H<G$ be a maximal subgroup. If $H$ is solvable, then the derived length of $H$ is at most $10$. The constant $10$ is realized only by the solvable maximal subgroups of the Fischer group $Fi_{{}_{23}}$.
Despite Theorem 1.3, it is easy to exhibit primitive nonsolvable groups of product type with solvable stabilizer of arbitrarily large derived length. Dealing with the product type case, we are able to answer in the affirmative the question at the end of [LS24]:
Theorem 1.4.
Let $G$ be a primitive group with solvable stabilizer. Then there exist two points whose pointwise stabilizer has derived length bounded by an absolute constant.
The proof of Theorem 1.4 is an application of Corollary 1.2, and it is perhaps surprising that, to answer a question about pointwise stabilizers of primitive nonsolvable groups, one has to deal with setwise stabilizers of arbitrary solvable permutation groups. The idea, usually applied in the context of imprimitive linear groups, is using a partition to construct a vector with small centralizer (see [MW04, Sec. 4]).
The paper is organized as follows. In Section 2 we study setswise stabilizers, proving Theorem 1.1 and other related results (we also answer [Bab22, Question 7c], and give an elementary proof of a theorem of Cameron, Neumann and Saxl). In Section 3, we prove Theorems 1.3 and 1.4. The methods are a mixture of group theory and combinatorics.
Acknowledgments: I am grateful to S. Eberhard for many useful conversations. After the public announcement of a first version of this paper, I received precious comments from D. Gluck, H.Y. Huang and S.V. Skresanov. The first asked for a result in the spirit of Corollary 2.11. The second noticed that the constant $10$ in Theorem 1.3 is attained by $Fi_{{}_{23}}$. The third pointed out Babai’s paper [Bab22], and helped me with the computer calculations in Lemma 2.6. Finally, I am thankful to L. Pyber for the many ideas he has given me in these years.
2. Setwise stabilizers of solvable groups
Let $G$ be a finite group. The derived series is defined by $G^{(0)}=G$, and
$G^{(n+1)}\>=\>[G^{(n)},G^{(n)}]$
for all $n\geq 0$. For a solvable group $G$, the derived length $dl(G)$ is the minimum $n$ such that $G^{(n)}=1$. For example, $dl(G)=1$ if and only if $G$ is abelian and $dl(G)\leq 2$ if and only if $G$ is metabelian.
When $G$ acts on a finite set $\Omega$, then $G$ acts naturally on the power set $\mathcal{P}(\Omega)$. If $\Delta\subseteq\Omega$, we write $C_{G}(\Delta)$ for the pointwise stabilizer, and $Stab_{G}(\Delta)$ for the setwise stabilizer. As usual, we say that $\Delta$ lies in a regular orbit if $Stab_{G}(\Delta)=1$. We say that $\Delta$ lies in a binary orbit if all the orbits of $Stab_{G}(\Delta)$ on $\Omega$ have length at most $2$, in a ternary orbit if all the orbits of $Stab_{G}(\Delta)$ on $\Omega$ have length at most $3$, and in a good orbit if all the orbits of $Stab_{G}(\Delta)$ on $\Omega$ have length at most $6$. Of course,
[…]
Remark 2.1.
The total number of orbits of $G$ on $\mathcal{P}(\Omega)$ is at least $|\Omega|+1$.
We now make a simple consideration. If $G\leqslant\mathop{\mathrm{Sym}}(\Omega)$ and $\Delta=\{x_{1},\ldots,x_{b}\}$ is a base (i.e., $C_{G}(\Delta)=1$), then
$Stab_{G}(\Delta)\cong Stab_{G}(\Delta)/C_{G}(\Delta)\leqslant\mathop{\mathrm{Sym}}(b).$
So, if $Stab_{G}(\Delta)$ is solvable, then $dl(Stab_{G}(\Delta))\leq f(b)$, where
$f(b)=\>\max\,\{dl(S):S\leqslant\mathop{\mathrm{Sym}}(b),S\mbox{ solvable}\}.$
We point out that $f(b)=O(\log b)$ [Dix68]. When combined with [Ser96], this gives that in a primitive solvable group there exists a setwise stabilizer of derived length at most $f(4)=3$. We are going to see that a bit more is true.
2.1. Primitive solvable groups
Gluck [Glu83] proved that, with finitely many exceptions, a primitive solvable group has a regular orbit on the power set. His result has been refined several times, for example in [Ser97, MW04]. The following version is tailored for attacking Theorem 1.1, and its proof is mostly computational.
Lemma 2.2.
Let $G\leqslant\mathop{\mathrm{Sym}}(\Omega)$ be a primitive solvable group. If $|\Omega|\geq 10$, then $G$ has at least $8$ regular orbits on $\mathcal{P}(\Omega)$. If $G$ has less than $5$ regular orbits on $\mathcal{P}(\Omega)$, then $G$ is one among the groups in Table 1.
Proof.
The first part follows from [MW04, Lemma 4.1(ii)]. The remaining cases have been handled with the library PrimitiveGroups in [GAP4]. ∎
Similarly, we remark the following:
Lemma 2.3.
Let $G\leqslant\mathop{\mathrm{Sym}}(\Omega)$ be a primitive solvable group. The following hold:
  • There exists a subset of $\Omega$ whose setwise stabilizer has all orbits on $\Omega$ of length at most $3$;
  • There exists a subset of $\Omega$ whose setwise stabilizer is metabelian. If $G\neq\mathop{\mathrm{AGL}}_{2}(3)$, then there exists a subset whose setwise stabilizer is abelian.
Proof.
As before, using [MW04, Lemma 4.1(ii)]. ∎
$\dfrac{1^{1}}{1_{1_{1}}}$ GAP Id Regular Orbits on $\mathcal{P}(\Omega)$ Binary Orbits on $\mathcal{P}(\Omega)$ Ternary Orbits on $\mathcal{P}(\Omega)$ Good Orbits on $\mathcal{P}(\Omega)$
$\ell_{1,2}$ $\ell_{2,2}$ $\ell_{3,2}$ $\ell_{6,2}$
$\dfrac{1^{1}}{1_{1_{1}}}$ $[2,1]$ $1$ $3$ $3$ $3$
$\dfrac{1^{1}}{1_{1_{1}}}$ $[3,1]$ $2$ $2$ $4$ $4$
$\dfrac{1^{1}}{1_{1_{1}}}$ $[3,2]$ 00 $2$ $4$ $4$
$\dfrac{1^{1}}{1_{1_{1}}}$ $[4,1]$ 00 $1$ $3$ $5$
$\dfrac{1^{1}}{1_{1_{1}}}$ $[4,2]$ 00 $1$ $3$ $5$
$\dfrac{1^{1}}{1_{1_{1}}}$ $[5,2]$ 00 $6$ $6$ $8$
$\dfrac{1^{1}}{1_{1_{1}}}$ $[5,3]$ 00 $2$ $2$ $6$
$\dfrac{1^{1}}{1_{1_{1}}}$ $[7,2]$ $2$ $16$ $16$ $16$
$\dfrac{1^{1}}{1_{1_{1}}}$ $[7,3]$ $4$ $4$ $10$ $10$
$\dfrac{1^{1}}{1_{1_{1}}}$ $[7,4]$ 00 $4$ $6$ $8$
$\dfrac{1^{1}}{1_{1_{1}}}$ $[8,1]$ $3$ $5$ $5$ $6$
$\dfrac{1^{1}}{1_{1_{1}}}$ $[8,2]$ 00 00 $3$ $6$
$\dfrac{1^{1}}{1_{1_{1}}}$ $[9,2]$ 00 $10$ $10$ $24$
$\dfrac{1^{1}}{1_{1_{1}}}$ $[9,3]$ $4$ $6$ $6$ $14$
$\dfrac{1^{1}}{1_{1_{1}}}$ $[9,4]$ $4$ $8$ $8$ $12$
$\dfrac{1^{1}}{1_{1_{1}}}$ $[9,5]$ 00 $4$ $4$ $12$
$\dfrac{1^{1}}{1_{1_{1}}}$ $[9,6]$ 00 00 $4$ $10$
$\dfrac{1^{1}}{1_{1_{1}}}$ $[9,7]$ 00 00 $4$ $10$
Table 1. Primitive solvable groups with less than $5$ regular orbits on $\mathcal{P}(\Omega)$.
As an example to help reading Table 1, we remark that PrimitiveGroup(4,2) is $\mathop{\mathrm{Sym}}(4)$. There is no regular orbit of $\mathop{\mathrm{Sym}}(4)$ on $\mathcal{P}(\{1,2,3,4\})$, the stabilizer of $\{1,2\}$ has two orbits of length $2$, and the stabilizers of $\{1\}$ and $\{1,2,3\}$ have orbits of length $3$. We also remark that PrimitiveGroup(8,2) is $\mathrm{A\Gamma L}_{1}(8)$, while PrimitiveGroup(9,7) is $\mathop{\mathrm{AGL}}_{2}(3)$.
We now report the following sharp result:
Lemma 2.4.
If $G\leqslant\mathop{\mathrm{Sym}}(\Omega)$ is a primitive solvable group, then there exist disjoint subsets $\Delta_{1},\Delta_{2},\Delta_{3}$ such that
$\cap_{j=1}^{3}Stab_{G}(\Delta_{j})=1.$
Proof.
See [MW04, Lemma 4.1(iii)]. ∎
2.2. Imprimitive groups, colorings
The standard argument to deal with imprimitive groups is [Glu83, Theorem 2], which was used by Gluck to show that permutation groups of odd order have a regular orbit on the power set. This does not work with a property such as “have all orbits of length at most $C$”, where $C$ is a constant. We use a different method based on good orbits and partitions that we are going to explain.
The distinguishing number of a permutation group $G$ on $\Omega$ is the smallest number of parts in a partition of $\Omega$ with the property that only the identity fixes every part. For example, this is $2$ if and only if $G$ has a regular orbit on the power set. For more details, see [BC11, Sec. 2.2]. We refine the notion as follows:
Definition.
Let $G\leqslant\mathop{\mathrm{Sym}}(\Omega)$, and $k\geq 2$. A $k$-coloring is a coloring of $\Omega$ with at most $k$ colors $\{1,\ldots,k\}$. A regular $k$-coloring is a $k$-coloring such that the identity is the only element of $G$ that fixes every color. A binary $k$-coloring is a $k$-coloring such that the stabilizer of the coloring in $G$ has all orbits of length at most $2$. Ternary $k$-colorings and good $k$-colorings are defined accordingly.
The total number of $k$-colorings of $\Omega$ is $k^{|\Omega|}$, but many of these are equivalent under the action of $G$. Fixed $G\leqslant\mathop{\mathrm{Sym}}(\Omega)$, for each $k\geq 2$ we write
[…]
for the numbers of $G$-orbits of regular, binary, ternary, and good $k$-colorings. It is obvious that, for every $i\leq j$ and $k$,
(2.1) $\ell_{i,k+1}\geq\ell_{i,k},\hskip 28.45274pt\mbox{ and }\hskip 28.45274pt\ell_{j,k}\geq\ell_{i,k}.$
We observe that describing a $2$-coloring is equivalent to describing a subset. Therefore, $\ell_{1,2}$, $\ell_{2,2}$, $\ell_{3,2}$ and $\ell_{6,2}$ coincide respectively with the numbers of regular, binary, ternary and good orbits on the power set. Restated in terms of colorings, Lemma 2.4 says that $\ell_{1,4}\geq 1$ for every primitive solvable group (one color is needed for the complement).
A portion of the proof of Theorem 1.1 relies on showing that $\ell_{i,6-i}\geq 5$ for all primitive solvable groups and each $i=1,2,3$. The following combinatorial lemma is useful to estimate $\ell_{i,k}$ for large $k$.
Lemma 2.5.
Fix $i\in\{1,2,3,6\}$. If $\ell_{i,2}\geq 1$, then $\ell_{i,k}\geq{k\choose 2}$. If $\ell_{i,j}>\ell_{i,j-1}$ for some $3\leq j\leq k$, then $\ell_{i,k}\geq{k\choose j}$.
Proof.
The inequality $\ell_{i,j}>\ell_{i,j-1}$ means that there exists a suitable coloring that involves precisely $j$ colors. Now suppose we have such a $j$-coloring. For each subset of cardinality $j$ of the $k$ available colors, we obtain an inequivalent suitable $j$-coloring. ∎
The following result requires a slightly deeper investigation. In the notation $\ell_{i,k}$, we write $\ell_{i,k}(G)$ to specify the group we are referring to.
Lemma 2.6.
Let $A$ be a primitive solvable group. Then $\ell_{2,3}(A)\geq 5$.
Proof.
Let $\Omega$ be the base set. By Lemma 2.2 and (2.1), we can assume $|\Omega|\leq 9$. Since we are interested in a lower bound on the number of orbits, it is enough to deal with the groups
[…]
It is easy to see that $\ell_{2,3}$ is at least $5$ in all these cases. In particular, using [GAP4], the command
OrbitsDomain( A , Tuples([1..k], NrMovedPoints(A)) , Permuted )
returns the orbits of $A$ on the $k$-colorings of $\Omega$. ∎
We prove Theorem 1.1 through the following stronger statement, that lends itself to an inductive argument.
Theorem 2.7.
Let $G\leqslant\mathop{\mathrm{Sym}}(\Omega)$ be a solvable permutation group. If $|\Omega|\geq 2$, then there are at least $3$ good orbits on $\mathcal{P}(\Omega)$. If there are less than $5$ good orbits on $\mathcal{P}(\Omega)$, then $|\Omega|\leq 3$.
Proof.
We work by induction on the cardinality of $\Omega$. Suppose that $\Omega_{1},\ldots,\Omega_{t}$ are the orbits of $G$ on $\Omega,$ and that we can find subsets $\Delta_{j}\subseteq\Omega_{j}$ such that $Stab_{G}(\Delta_{j})$ has orbits of length at most $6$ on $\Omega_{j}$, for each $j=1,\ldots,t$. Then $Stab_{G}(\sqcup_{j}\>\Delta_{j})$ will surely work. It follows that we can suppose that $G$ is transitive. Primitive groups can be handled with Lemma 2.2 and Table 1. Let $\Omega=\Omega_{1}\times\Omega_{2}$, with $|\Omega_{1}|=a\geq 2$. Let $G\leqslant B\wr_{a}A$, with $A$ primitive on $\Omega_{1}$, and $B$ transitive on $\Omega_{2}$. By the inductive hypothesis applied to $\Omega_{2}$, we are in one of the following cases:
  • $B$ has at least $5$ good orbits on $\mathcal{P}(\Omega_{2})$.
    We are going to use the good orbits of $B$ on $\mathcal{P}(\Omega_{2})$, say $\Sigma_{1}^{B},\ldots,\Sigma_{5}^{B}$, as colors for $\Omega_{1}$. Indeed, for each regular $5$-coloring of $\Omega_{1}$, say $\Omega_{1}=\sqcup_{j=1}^{5}\Delta_{j}$, we can construct a good orbit $\Delta^{G}$ for $G$ as follows:
    $\Delta=\{(x,y)\in\Omega_{1}\times\Omega_{2}:x\in\Delta_{j}\mbox{ and }y\in\Sigma_{j}\mbox{ for some fixed }j\}.$
    The setwise stabilizer of $\Delta$ in $G$ can be embedded in the semidirect product between the direct product of the stabilizers of the $\Sigma_{j}$’s in $B$ (which have orbits of length at most $6$), and the stabilizer of the regular $5$-coloring in $A$ (which is trivial). Inequivalent $5$-colorings of $\Omega_{1}$ provide inequivalent good orbits of $\Omega$. In particular, there are at least $\ell_{1,5}(A)$ good orbits of $G$. Now $\ell_{1,4}(A)\geq 1$ by Lemma 2.4, and so, from Lemma 2.5,
    $\ell_{1,5}(A)\geq\min_{j=2,3,4}{5\choose j}=5.$
  • $B\cong\mathop{\mathrm{Sym}}(2)$.
    Here $\ell_{2,2}(B)=3$. Using the binary orbits of $B$, and the ternary $3$-colorings of $\Omega_{1}$, as before we can construct $\ell_{3,3}(A)$ good orbits of $G$. In fact, the length of any orbit will be bounded by $2\cdot 3=6$. Now $\ell_{3,3}(A)\geq\ell_{2,3}(A)\geq 5$ by (2.1) and Lemma 2.6.
  • $B\leqslant\mathop{\mathrm{Sym}}(3)$.
    We can assume $B\cong\mathop{\mathrm{Sym}}(3)$, so that $\ell_{3,2}(B)=4$. Using the ternary orbits of $B$, and the binary $4$-colorings of $\Omega_{1}$, as before we can construct $\ell_{2,4}(A)$ good orbits of $G$. In fact, the length of any orbit will be bounded by $3\cdot 2=6$. Now $\ell_{2,4}(A)\geq\ell_{2,3}(A)\geq 5$ by (2.1) and Lemma 2.6. ∎
The cardinality of $\Delta$ in Theorem 1.1 grows linearly with the cardinality of $\Omega$, which is necessary as shown by $B\wr K$ when $B$ has some orbit of length exceeding $6$. The bound on the derived length follows easily:
Proof of Corollary 1.2.
Let $\Delta\subseteq\Omega$ be as provided by Theorem 1.1. Then $Stab_{G}(\Delta)$ can be embedded in a direct product of copies of $\mathop{\mathrm{Sym}}(6)$. Every solvable subgroup of $\mathop{\mathrm{Sym}}(6)$ has derived length at most $3$, and the proof follows. ∎
The next example provides a transitive solvable group where no subset gives a metabelian stabilizer.
Example 2.8.
Let $B=\mathop{\mathrm{Sym}}(4)\wr_{4}\mathop{\mathrm{Sym}}(4)$ act imprimitively on $\Omega=\{a,b,c,d\}\times\{1,2,3,4\}$. It is not hard to see that $B$ has a unique orbit on the power set whose setwise stabilizer is metabelian, namely $\Delta^{B}$, where
$\Delta\>=\>\{(a,1),(b,1),(b,2),(b,3),(c,1),(c,2),(d,1),(d,2)\}.$
The setwise stabilizer $H$ of $\Delta$ in $B$ is isomorphic to $\mathop{\mathrm{Sym}}(3)\times\mathop{\mathrm{Sym}}(3)\times D_{8}$. Now let $G=B\wr_{2}\mathop{\mathrm{Sym}}(2)$ act on $\Omega\times\{\alpha,\beta\}$. If we color both $\Omega\times\{\alpha\}$ and $\Omega\times\{\beta\}$ using $\Delta$, then the resulting setwise stabilizer will be isomorphic to $H\wr_{2}\mathop{\mathrm{Sym}}(2)$, whose derived length is $3$. Otherwise, we are forced to color a subset, say of $\Omega\times\{\alpha\}$, which does not provide a metabelian setwise stabilizer in $B$. So there is no hope of getting a metabelian setwise stabilizer in $G$.
The next similar example involves $\mathrm{A\Gamma L}_{1}(8)$, and provides a transitive solvable group where no subset gives a stabilizer with all orbits of length at most $5$.
Example 2.9.
We first notice that $\mathrm{A\Gamma L}_{1}(8)$ has precisely $4$ orbits on the power set whose stabilizers have orbits of length at most $4$. Let $B=\mathrm{A\Gamma L}_{1}(8)\wr_{4}\mathop{\mathrm{Sym}}(4)$ act imprimitively on a set $\Omega$ of cardinality $32$. It is not hard to see that $B$ has a unique orbit $\Delta^{B}$ on the power set whose setwise stabilizer has orbits of length at most $5$ (actually $4$). Now let $G=B\wr_{2}\mathop{\mathrm{Sym}}(2)$ act on $\Omega\times\{\alpha,\beta\}$. If we color both $\Omega\times\{\alpha\}$ and $\Omega\times\{\beta\}$ using $\Delta$, then the resulting setwise stabilizer will have an orbit of length $4\cdot 2=8$. Otherwise, we are forced to color a subset, say of $\Omega\times\{\alpha\}$, which does not provide a setwise stabilizer (in $B$) with all orbits of length at most $5$. So there is no hope of getting such a stabilizer in $G$.
2.3. A result about $3$-colorings
With the same ideas used in the proof of Theorem 2.7, it is possible to obtain the following result concerning $3$-colorings. We observe that Corollary 2.11 answers another question of Babai [Bab22, Question 7c].
Theorem 2.10.
Let $G$ be a solvable permutation group on the finite set $\Omega$. Then there exist at least $5$ orbits of $3$-colorings of $\Omega$ whose stabilizers have all orbits of length at most $2$.
Proof.
We have to prove that $\ell_{2,3}(G)\geq 5$. We can suppose that $G$ is transitive. Primitive groups can be handled with Lemma 2.6. Let $\Omega=\Omega_{1}\times\Omega_{2}$, with $|\Omega_{1}|=a\geq 2$. Let $G\leqslant B\wr_{a}A$, with $A$ primitive on $\Omega_{1}$, and $B$ transitive on $\Omega_{2}$. By the inductive hypothesis, we have that $\ell_{2,3}(B)\geq 5$. Arguing as in the proof of Theorem 2.7, using these binary $3$-colorings of $\Omega_{2}$ as colors for $\Omega_{1}$, and the regular $5$-colorings of $\Omega_{1}$, we can construct $\ell_{1,5}(A)$ binary $3$-colorings of $\Omega$. Finally, from Lemmas 2.4 and 2.5, we have $\ell_{1,5}(A)\geq 5$. ∎
Corollary 2.11.
Let $G$ be a solvable permutation group on the finite set $\Omega$. Then there exists a $3$-coloring of $\Omega$ whose stabilizer is an elementary abelian $2$-group.
As already noticed by Babai, not even a generic $2$-group can be obtained with just two colors: take $\mathop{\mathrm{Sym}}(4)\wr K$ for any solvable $K$ that is not a $2$-group.
2.4. A CFSG-free proof for arbitrary primitive groups
The foundation stone to prove a result like Theorem 1.1 is the fact that a primitive solvable group of large degree has a regular orbit on the power set [Glu83]. With the trivial exceptions of the alternating and symmetric groups, the same is actually true in the nonsolvable case, but the known proofs of Cameron, Neumann, Saxl and Seress [CNS84, Ser97] require the classification of the finite simple groups. We observe that a more recent elementary result of Sun and Wilmes [SW15, SW16] has the following consequence.
Theorem 2.12.
Let $G$ be a primitive group not containing $\mathop{\mathrm{Alt}}(n)$. With only finitely many exceptions, $G$ has a regular orbit on the power set.
Proof.
Let $n$ be the degree of $G$. We mimic the short proof in [CNS84], which uses the notions of base size and minimal degree in a clever way. At the crucial point, they obtain $|G|\geq 2^{\sqrt{n/2}}$, and use the classification to remain only with Johnson and Hamming groups, which are handled separetely. While Babai’s theorem [Bab81] is not enough, [SW16, Corollary 1.6] is sufficient for this job, having a thresold of $2^{n^{1/3}(\log n)^{7/3}}$. ∎
Let $G$ be an imprimitive group. To glue together the various pieces in Theorem 2.12, one has to be careful that large alternating groups are not induced on any subset. This fact makes the possible statements more complicated. One consequence is the following:
Theorem 2.13.
Let $G$ be a finite permutation group on $\Omega$ which does not involve $\mathop{\mathrm{Alt}}(n)$ for any $n\geq n_{0}$. Then there is a subset of $\Omega$ whose setwise stabilizer has all orbits of length bounded by a function of $n_{0}$.
3. Primitive groups with solvable stabilizer
We would like to remark that, the idea of fixing a bounded number of points to obtain something interesting, is completely hopeless in imprimitive groups. This is a key difference with respect to setwise stabilizers (see Corollaries 1.2 and 2.11, for example).
Example 3.1.
Let $B$ be a transitive group with stabilizer $B_{0}$, and let $C_{k}$ be the cyclic group on $k$ points. Let $G=B\wr_{k}C_{k}$ act imprimitively on $|B:B_{0}|k$ points. The stabilizer of a point in $G$ is $H\cong B_{0}\times B^{k-1}$. For each $b\geq 1$, the intersection of any $b$ conjugates of $H$ will contain a copy of $B^{k-b}$, showing that fixing less than $k$ points is useless.
Ultimately, the goal of this section is to prove Theorem 1.4. In the case of primitive solvable groups, this is [LS24, Theorem 1.1]. The proof of this part is really about finding a good centralizer in every irreducible solvable linear group, which was done by Yang [Yan09] using ideas in [MW04]. In general, as explained by Li and Zhang in [LZ11, Theorem 1.1], a primitive nonsolvable group with solvable stabilizer is almost simple or of a specific product type.
3.1. Almost simple groups
Theorem 1.3 will follow from the tables in [LZ11, Sec. 10], together with a few properties of finite simple groups and linear groups. We report a strong version of the Schreier conjecture, which says that the group $\mathop{\mathrm{Out}}(T)$ of the outer automorphisms of a finite simple group is solvable.
Lemma 3.2.
If $T$ is a finite simple group, then $\mathop{\mathrm{Out}}(T)$ is solvable of derived length at most $3$. More precisely:
  • (i)
    If $T$ is cyclic, alternating, or a sporadic simple group, then $\mathop{\mathrm{Out}}(T)$ is abelian;
  • (ii)
    $dl(\mathop{\mathrm{Out}}(T))=3$ if and only if $T$ is the orthogonal group $\mathrm{P}\Omega^{+}_{8}(q)$ for $q$ odd.
Proof.
This follows from the classification of the finite simple groups. In particular, see [CCN+85, Pag. xvi, Table 5]. For finite simple groups of Lie Type, $\mathop{\mathrm{Out}}(T)$ can be described as an extension of three groups: diagonal, field and graph automorphisms (the field and graph automorphisms commute). These groups are abelian except for the graph automorphisms group of $\mathrm{P}\Omega^{+}_{8}(q)$, which is isomorphic to $\mathop{\mathrm{Sym}}(3)$. ∎
Remark 3.3.
In an almost simple group $G$, the socle $T$ is the unique minimal normal subgroup. In fact, let $N\neq T$ be another minimal normal subgroup. Then $[N,T]=1$ and so $N\subseteq C_{G}(T)=1$. This implies that the solvable radical of $G$ is trivial and that every solvable subgroup of $G$ is core-free.
Proof of Theorem 1.3.
Let $T$ be the socle of $G$, and let $G_{0}$ be the normal subgroup containing $T$ which is described in [LZ11, Pag. 4]. We have that $H$ is an extension of $H_{0}=H\cap G_{0}$ with
$H/H_{0}\cong G_{0}H/G_{0}\cong G/G_{0},$
which is a section of $G/T$ and so of $\mathop{\mathrm{Out}}(T)$. It follows that
$dl(H)\leq dl(H_{0})+dl(G/G_{0}).$
Looking at the Tables 14-20 in [LZ11], we observe that one of the following holds:
  • (i)
    $G_{0}$, and so $H_{0}$, is one of finitely many groups;
  • (ii)
    $G_{0}$ is symmetric or alternating, and $H_{0}$ has derived length bounded by $6$;
  • (iii)
    $G_{0}$ is a simple group of Lie type of bounded rank (recall that exceptional groups lie in this class);
  • (iv)
    $G_{0}$ is $\mathop{\mathrm{PSL}}_{r}(q)$ or $\mathop{\mathrm{PSU}}_{r}(q)$, and $H_{0}$ is metabelian.
To deal with (iii) uniformly, we use the fact that a solvable linear group has derived length bounded by a function on the rank [Dix68]. The (relatively high) constant $10$ is attained by the Fischer group $Fi_{{}_{23}}$, which makes smooth the remaining computations. We point out some mistakes in [LZ11, Table 15], and rely instead on the [GAP4] library CTblLib (v1.3.9). [^2] One mistake is in correspondence of the solvable maximal subgroups of $Fi_{{}_{23}}$. For example, a potential obstruction is given by a certain conjugacy class of solvable maximal subgroups $H$ of the baby monster, which seems to be not available on a computer. With the ATLAS notation [CCN+85], $H$ is isomorphic to $3^{11}.(S4\times 2S4)$. A group of order $3^{11}$ has nilpotency class at most $10$ and derived length at most $4$, showing that $dl(H)\leq 8$ in this case. ∎
A weaker version of Theorem 1.3 says that there exists a constant $C$ such that if $G$ is a finite group with a solvable maximal subgroup $H$, and the derived length of $H$ is more than $C$, then $G$ is not simple. It would be interesting to find a proof of this fact which does not rely on the classification of the finite simple groups.
3.2. Groups of product type
Let $X$ be a primitive almost simple group with solvable point stabilizer $Y<X$, and let $T\lhd X$ be its socle. Let $d\geq 2$, and let $S\leqslant\mathop{\mathrm{Sym}}(d)$ be a transitive solvable group. By [LZ11, Theorem 1.1], the primitive group $G$ that we care about satisfies
$T\wr_{d}S\>\leqslant\>G\>\leqslant\>X\wr_{d}S.$
(See also [Bur21, Sec. 8.2].) Let $V\cong X^{d}$ be the base subgroup of $X\wr_{d}S$, and let $H<G$ be the stabilizer of a point in the product action. Let $W\cong Y^{d}$ be the subgroup of $V$ such that $H=WS\cap G$. Stabilizing two points corresponds to intersecting $H$ with one of its conjugates. To prove Theorem 1.4, we will find $v\in G\cap V$ such that the derived length of
$I_{v}\>=\>WS\cap(WS)^{v}$
is bounded. The proof will follow, because $H\cap H^{v}\leqslant I_{v}$, and the derived length of a subgroup is bounded by the derived length of the group.
Proof of Theorem 1.4.
Let $\pi\colon G\to S$ be the natural projection, and let $v\in V$. Since $Ker(\pi)=G\cap V$, we have
$dl(I_{v})\leq dl(I_{v}\cap V)+dl(\pi(I_{v})).$
We observe that
$I_{v}\cap V\subseteq WS\cap V=W.$
This gives $I_{v}\cap V=I_{v}\cap W$, and therefore we can write
$dl(I_{v}\cap V)\leq dl(W)=dl(Y).$
Now $dl(Y)$ is bounded by Theorem 1.3, and the rest of the proof deals with $\pi(I_{v})$. It is not hard to see that $C_{S}(v)=S\cap S^{v}\subseteq\pi(I_{v})$, but in general $\pi(I_{v})$ can be quite bigger. For every $w\in W$, $\sigma\in S$ and $v\in V$, an explicit computation gives
$(w,\sigma)^{v}=(v^{-1}wv^{\sigma},\sigma).$
This implies that
$\pi(I_{v})\>=\>\{\sigma\in S:\>W\cap v^{-1}Wv^{\sigma}\neq\emptyset\}.$
Now let $v=(x_{1},\ldots,x_{d})$. We have
$\pi(I_{(x_{1},\ldots,x_{d})})=\{\sigma\in S:\>Y\cap x_{i}^{-1}Yx_{\sigma(i)}\neq\emptyset\mbox{ for each }i=1,\ldots,d\}.$
Since
$Y\cap x_{i}^{-1}Yx_{\sigma(i)}\neq\emptyset\hskip 28.45274pt\mbox{ if and onlyif }\hskip 28.45274ptx_{\sigma(i)}\in Yx_{i}Y,$
we obtain
(3.1) $\pi(I_{v})=\{\sigma\in S:\>x_{\sigma(i)}\in Yx_{i}Y\mbox{ for each }i=1,\ldots,d\}.$
We are going to look at the double cosets of $Y$ as colors. Let $c_{1},\ldots,c_{r}\in X$ be representatives of these double cosets. For any $v\in V$, and $j\in\{1,\ldots,r\}$, let
$\Delta_{j}=\{i\in\{1,\ldots,d\}:x_{i}\in Yc_{j}Y\}.$
Of course the $\Delta_{j}$’s form a partition of $\{1,\ldots,d\}$. The equality in (3.1) shows that $\pi(I_{v})$ only depends on this partition, and in particular
(3.2) $\pi(I_{v})=\cap_{j=1}^{r}Stab_{S}(\Delta_{j})=\cap_{j=1}^{r-1}Stab_{S}(\Delta_{j}).$
To construct $v$, we will use only two elements $c_{1}$ and $c_{2}$ (and so two sets $\Delta_{1}$ and $\Delta_{2}=\{1,\ldots,d\}\setminus\Delta_{1}$). One reason is that we want $v$ to lie in $G\cap V$, and not just in $V$. This is assured if we choose representatives $c_{i}$’s which are in $T$. We can surely choose $c_{1}=1$ and $c_{2}\in T\setminus\{1\}$, because otherwise $T\leqslant Y$, which is impossible because $Y$ is solvable. Finally, choosing $\Delta_{1}=\Delta\subseteq\{1,\ldots,d\}$ as provided by Corollary 1.2, from (3.2) we get $dl(\pi(I_{v}))\leq 3$, concluding $dl(I_{v})\leq 13$ as desired. ∎
Remark 3.4.
In the proof of Theorem 1.4, it might seem wasteful to construct $v$ with only two elements $c_{1}$ and $c_{2}$. As explained, one reason is that we want $v\in G$, but there is also a different motivation. In fact, it is well known that the number $r$ of double cosets of $Y$ in $X$ equals the rank of the primitive group $X$ with solvable stabilizer $Y$, and there actually exist many doubly transitive groups of this type (as $\mathop{\mathrm{Sym}}(5)$, for example). A more detailed analysis of highly transitive almost simple groups, and the additional use of Corollary 2.11, might help to find the right constant in Theorem 1.4. We remark that this constant is not even known in the case of solvable groups, where it lies somewhere between $4$ and $9$ (see [LS24, Remark 2.9]).
Corollary 3.5.
Let $\delta\in\{2,3,4\}$. Let $G\leqslant\mathop{\mathrm{Sym}}(\Omega)$ be a primitive group with solvable stabilizer. Then there exists $\Delta\subseteq\Omega$, $|\Delta|=\delta$, such that $Stab_{G}(\Delta)$ is solvable of derived length bounded by an absolute constant.
Proof.
Let $\{x_{1},x_{2}\}\subseteq\Omega$ be as provided by Theorem 1.4. If $\{x_{1},x_{2}\}\subseteq\Delta$, then $C_{G}(\Delta)$ is solvable of bounded derived length. If $|\Delta|\leq 4$, then $Stab_{G}(\Delta)/C_{G}(\Delta)\leqslant\mathop{\mathrm{Sym}}(4)$, and it follows that $Stab_{G}(\Delta)$ is solvable of bounded derived lenght. ∎

### v3

@@
 We will see that this can be done in a few interesting situations.
-Our first main result is the following:
+Regarding subsets, Babai [Bab22] has proved that in every solvable permutation group there exists a set-stabilizer whose orbits have length bounded by a nonexplicit absolute constant.
+Answering [Bab22, Question 7a], our first main result provides the best possible constant:
 Theorem 1.1.
-Let $G\leqslant\mathop{\mathrm{Sym}}(\Omega)$ be a solvable permutation group. Then there exists a subset of $\Omega$ whose setwise stabilizer has all orbits of length at most $6$.
+Let $G\leqslant\mathop{\mathrm{Sym}}(\Omega)$ be a solvable permutation group.
+Then there exists a subset of $\Omega$ whose setwise stabilizer has all orbits of length at most $6$.
+The constant $6$ is the best possible.
 The following consequence is immediate.
@@
 Let $G\leqslant\mathop{\mathrm{Sym}}(\Omega)$ be a solvable permutation group. Then there exists a subset of $\Omega$ whose setwise stabilizer has derived length at most $3$.
-For primitive solvable groups, we can find a subset whose stabilizer has orbits of length at most $3$ and is metabelian.
-In the beginning we hoped to prove the same for all transitive solvable groups, but this happened to be false (see Examples 2.8 and 2.9).
-In fact, the results above are the best possible, and answer two questions of Babai [Bab22, Questions 7a-7b].
-Theorem 1.1 seems to be the first statement concerning stabilizers of imprimitive groups where the property involved is not preserved by group extensions, and part of the difficulty of the proof is in finding the right statement to prove by induction.
-[^1] Babai [Bab22, Theorem 12.1] proves the same result with a nonexplicit constant.
+It happens that metabelian stabilizers need not to exist in general (see Example 2.11), so Corollary 1.2 answers [Bab22, Question 7b].
+Incidentally, the same example answers negatively a question in [Glu24] on the existence of supersolvable set-stabilizers.
+We stress that the two constants $6$ and $3$ above turned out to be a little higher than what initially expected, and better results can be obtained for primitive solvable groups.
+As far as we know, Theorem 1.1 is the first statement concerning stabilizers of imprimitive groups where the property involved is not preserved by a group extension, and part of the difficulty of the proof is in finding the right statement to prove by induction.
+With similar ideas, we are able to show that, in every solvable permutation group, stabilizing two disjoint subsets can force the orbits to have length at most $2$.
+This gives a strong positive answer to [Bab22, Question 7c].
+Moreover, we give some elementary proofs concerning nonsolvable groups: we provide an easier proof of a theorem of Cameron [Cam86], and answer [Bab22, Question 6a].
+In particular, we prove the following:
+Theorem 1.3.
+Let $G$ be any permutation group on a finite set. Then there exists a setwise stabilizer whose nonalternating composition factors have order bounded by an absolute constant.
+This is new even using the Classification of Finite Simple Groups.
 We come to the question of fixing two points.
 Extending the work of Seress [Ser96], Burness [Bur21] showed that in a primitive group with solvable stabilizer, there always exist $5$ points whose pointwise stabilizer is trivial.
+The hypothesis of having a solvable stabilizer is natural, if we aim to structural results on the stabilizer of two points.
 For primitive solvable groups, in [LS24] the authors observed that there exist two points whose pointwise stabilizer has bounded derived length.
-On the other hand, motivated by various applications, Li and Zhang [LZ11] have classified the solvable maximal subgroups of the almost simple groups (see also [Bur21]).
+On the other hand, motivated by various applications, Li and Zhang [LZ11] have classified the solvable maximal subgroups of the almost simple groups.
 Combining their work with some properties of the finite simple groups, we remark the following:
@@
 If $H$ is solvable, then the derived length of $H$ is at most $10$.
-The constant $10$ is realized only by the solvable maximal subgroups of the Fischer group $Fi_{{}_{23}}$.
+The constant $10$ is attained only by the solvable maximal subgroups of the Fischer group $Fi_{{}_{23}}$.
 Despite Theorem 1.4, it is easy to exhibit primitive nonsolvable groups of product type with solvable stabilizer of arbitrarily large derived length.
@@
 Let $G$ be a primitive group with solvable stabilizer. Then there exist two points whose pointwise stabilizer has derived length bounded by an absolute constant.
-The proof of Theorem 1.4 is an application of Corollary 1.2, and it is perhaps surprising that, to answer a question about pointwise stabilizers of primitive nonsolvable groups, one has to deal with setwise stabilizers of arbitrary solvable permutation groups.
+The proof of Theorem 1.5 is an application of Corollary 1.2, and it is perhaps surprising that, to answer a question about pointwise stabilizers of primitive nonsolvable groups, one has to deal with set-stabilizers of arbitrary solvable permutation groups.
 The idea, usually applied in the context of imprimitive linear groups, is using a partition to construct a vector with small centralizer (see [MW04, Sec.
 4]).
+Our argument involves a double coset decomposition and a neat bound gives the constant $13$.
+It is likely that a more in-depth study of almost simple groups leads to an improvement, but the best possible constant is not even known for solvable groups [LS24].
 The paper is organized as follows.
-In Section 2 we study setswise stabilizers, proving Theorem 1.1 and other related results (we also answer [Bab22, Question 7c], and give an elementary proof of a theorem of Cameron, Neumann and Saxl).
-In Section 3, we prove Theorems 1.3 and 1.4.
+In Section 2 we study set-stabilizers, proving Theorem 1.1 and the other related results.
+In Section 3, we prove Theorems 1.4 and 1.5.
 The methods are a mixture of group theory and combinatorics.
@@
 Skresanov.
-The first asked for a result in the spirit of Corollary 2.11.
-The second noticed that the constant $10$ in Theorem 1.3 is attained by $Fi_{{}_{23}}$.
-The third pointed out Babai’s paper [Bab22], and helped me with the computer calculations in Lemma 2.6.
+The first asked for a structural result in the spirit of Corollary 2.14 (see also [Glu24]).
+The second noticed that the (relatively high) constant $10$ in Theorem 1.4 is attained by $Fi_{{}_{23}}$, and that the constant in Theorem 1.5 is at least $5$ for almost simple groups.
+The third brought the article [Bab22] to my attention, and helped me with the computer calculations in Lemma 2.7.
+After this, I did a substantial rewrite of the paper and had useful conversations with L.
+Babai.
 Finally, I am thankful to L.
 Pyber for the many ideas he has given me in these years.
-2. Setwise stabilizers of solvable groups
+2. Breaking symmetry with set-stabilizers
 Let $G$ be a finite group. The derived series is defined by $G^{(0)}=G$, and
@@
 If $\Delta\subseteq\Omega$, we write $C_{G}(\Delta)$ for the pointwise stabilizer, and $Stab_{G}(\Delta)$ for the setwise stabilizer.
-As usual, we say that $\Delta$ lies in a regular orbit if $Stab_{G}(\Delta)=1$.
-We say that $\Delta$ lies in a binary orbit if all the orbits of $Stab_{G}(\Delta)$ on $\Omega$ have length at most $2$, in a ternary orbit if all the orbits of $Stab_{G}(\Delta)$ on $\Omega$ have length at most $3$, and in a good orbit if all the orbits of $Stab_{G}(\Delta)$ on $\Omega$ have length at most $6$.
-Of course,
+For each $g\in G$, it is easy to see that the stabilizers of $\Delta$ and $\Delta^{g}$ are conjugate subgroups.
+As usual, we say that $\Delta$ lies in a regular orbit (i.e., the orbit $\Delta^{G}$) if $Stab_{G}(\Delta)=1$.
+Of course, as subgroups of $G$, the stabilizers still act on $\Omega$.
+Definition.
+For $i\geq 1$, we say that $\Delta\subseteq\Omega$ lies in a $i$-regular orbit if all the orbits of $Stab_{G}(\Delta)$ on $\Omega$ have length at most $i$.
+The $1$-regular orbits are precisely the genuine regular orbits.
+Moreover, unless explicitly stated otherwise, the $6$-regular orbits will also be called good orbits.
+(This is motivated by Theorem 1.1.) We note that, by definition,
 […]
@@
 We point out that $f(b)=O(\log b)$ [Dix68].
-When combined with [Ser96], this gives that in a primitive solvable group there exists a setwise stabilizer of derived length at most $f(4)=3$.
-We are going to see that a bit more is true.
+When combined with [Ser96], this gives that in every primitive solvable group there exists a set-stabilizer of derived length at most $f(4)=3$.
+We are going to see that a little more is true.
 2.1. Primitive solvable groups
@@
 The first part follows from [MW04, Lemma 4.1(ii)]. The remaining cases have been handled with the library PrimitiveGroups in [GAP4]. ∎
-Similarly, we remark the following:
-Lemma 2.3.
-Let $G\leqslant\mathop{\mathrm{Sym}}(\Omega)$ be a primitive solvable group. The following hold:
-  • There exists a subset of $\Omega$ whose setwise stabilizer has all orbits on $\Omega$ of length at most $3$;
-  • There exists a subset of $\Omega$ whose setwise stabilizer is metabelian. If $G\neq\mathop{\mathrm{AGL}}_{2}(3)$, then there exists a subset whose setwise stabilizer is abelian.
-Proof.
-As before, using [MW04, Lemma 4.1(ii)]. ∎
-$\dfrac{1^{1}}{1_{1_{1}}}$ GAP Id Regular Orbits on $\mathcal{P}(\Omega)$ Binary Orbits on $\mathcal{P}(\Omega)$ Ternary Orbits on $\mathcal{P}(\Omega)$ Good Orbits on $\mathcal{P}(\Omega)$
-$\ell_{1,2}$ $\ell_{2,2}$ $\ell_{3,2}$ $\ell_{6,2}$
+$\dfrac{1^{1}}{1_{1_{1}}}$ GAP Id Regular orbits on $\mathcal{P}(\Omega)$ $2$-regular orbits on $\mathcal{P}(\Omega)$ $3$-regular orbits on $\mathcal{P}(\Omega)$ Good orbits on $\mathcal{P}(\Omega)$
+$\ell_{2,1}$ $\ell_{2,2}$ $\ell_{2,3}$ $\ell_{2,6}$
 $\dfrac{1^{1}}{1_{1_{1}}}$ $[2,1]$ $1$ $3$ $3$ $3$
@@
 There is no regular orbit of $\mathop{\mathrm{Sym}}(4)$ on $\mathcal{P}(\{1,2,3,4\})$, the stabilizer of $\{1,2\}$ has two orbits of length $2$, and the stabilizers of $\{1\}$ and $\{1,2,3\}$ have orbits of length $3$.
-We also remark that PrimitiveGroup(8,2) is $\mathrm{A\Gamma L}_{1}(8)$, while PrimitiveGroup(9,7) is $\mathop{\mathrm{AGL}}_{2}(3)$.
-We now report the following sharp result:
+We also note that PrimitiveGroup(8,2) is $\mathrm{A\Gamma L}_{1}(8)$, while PrimitiveGroup(9,7) is $\mathop{\mathrm{AGL}}_{2}(3)$.
+The following result is obtained with similar considerations:
+Lemma 2.3.
+Let $G\leqslant\mathop{\mathrm{Sym}}(\Omega)$ be a primitive solvable group. The following hold:
+  • There exists a subset of $\Omega$ whose setwise stabilizer has all orbits on $\Omega$ of length at most $3$;
+  • There exists a subset of $\Omega$ whose setwise stabilizer is metabelian. If $G\neq\mathop{\mathrm{AGL}}_{2}(3)$, then there exists a subset whose setwise stabilizer is abelian.
+Proof.
+As before, using [MW04, Lemma 4.1(ii)], and the library PrimitiveGroups in [GAP4].
+We observe that $\mathrm{A\Gamma L}_{1}(8)$ and PrimitiveGroup(9,6) do not have any $2$-regular orbit on $\mathcal{P}(\Omega)$, but still have some abelian set-stabilizer.
+∎
+We conclude this subsection with the following result:
 Lemma 2.4.
 If $G\leqslant\mathop{\mathrm{Sym}}(\Omega)$ is a primitive solvable group, then there exist disjoint subsets $\Delta_{1},\Delta_{2},\Delta_{3}$ such that
-$\cap_{j=1}^{3}Stab_{G}(\Delta_{j})=1.$
+[…]
 Proof.
-See [MW04, Lemma 4.1(iii)]. ∎
-2.2. Imprimitive groups, colorings
+This is [MW04, Lemma 4.1(iii)]. As before, the proof works by checking the finitely many cases with $|\Omega|\leq 9$. ∎
+Lemma 2.4 is sharp, as shown by $\mathop{\mathrm{Sym}}(4)$.
+2.2. Good colorings
 The standard argument to deal with imprimitive groups is [Glu83, Theorem 2], which was used by Gluck to show that permutation groups of odd order have a regular orbit on the power set.
@@
 We use a different method based on good orbits and partitions that we are going to explain.
-The distinguishing number of a permutation group $G$ on $\Omega$ is the smallest number of parts in a partition of $\Omega$ with the property that only the identity fixes every part.
-For example, this is $2$ if and only if $G$ has a regular orbit on the power set.
-For more details, see [BC11, Sec.
-2.2].
-We refine the notion as follows:
+A key observation, already in the lines of the proof of [Bab22, Theorem 12.1], is that studying stabilizers of partitions can also be useful to obtain results concerning stabilizers of a single subset.
+We need some more definitions.
+An asymmetric coloring of a permutation group $G$ on $\Omega$ is a partition of $\Omega$ with the property that only the identity fixes every part.
+We refer the reader to [Bab22, Sec.
+2] for some history concerning breaking-symmetry results and some conflicting terminology.
+We refine this notion as follows:
 Definition.
 Let $G\leqslant\mathop{\mathrm{Sym}}(\Omega)$, and $k\geq 2$.
-A $k$-coloring is a coloring of $\Omega$ with at most $k$ colors $\{1,\ldots,k\}$.
-A regular $k$-coloring is a $k$-coloring such that the identity is the only element of $G$ that fixes every color.
-A binary $k$-coloring is a $k$-coloring such that the stabilizer of the coloring in $G$ has all orbits of length at most $2$.
-Ternary $k$-colorings and good $k$-colorings are defined accordingly.
-The total number of $k$-colorings of $\Omega$ is $k^{|\Omega|}$, but many of these are equivalent under the action of $G$. Fixed $G\leqslant\mathop{\mathrm{Sym}}(\Omega)$, for each $k\geq 2$ we write
+A $k$-coloring is a coloring of $\Omega$ with at most $k$ colors.
+For $i\geq 1$, an $i$-asymmetric $k$-coloring is a $k$-coloring such that the stabilizer of the coloring in $G$ has all orbits on $\Omega$ of length at most $i$.
+The $1$-asymmetric colorings are precisely the genuine asymmetric colorings.
+Unless explicitly stated otherwise, the $6$-asymmetric colorings will also be called good colorings.
+The total number of $k$-colorings of $\Omega$ is $k^{|\Omega|}$, but many of these are equivalent under the action of $G$.
+Fixed $G\leqslant\mathop{\mathrm{Sym}}(\Omega)$, for each $k\geq 2$ we write
 […]
-for the numbers of $G$-orbits of regular, binary, ternary, and good $k$-colorings. It is obvious that, for every $i\leq j$ and $k$,
-(2.1) $\ell_{i,k+1}\geq\ell_{i,k},\hskip 28.45274pt\mbox{ and }\hskip 28.45274pt\ell_{j,k}\geq\ell_{i,k}.$
+for the numbers of $G$-orbits of asymmetric, $2$-asymmetric, $3$-asymmetric and good $k$-colorings.
+When necessary, we write $\ell_{k,i}(G)$ to stress the group $G$ we are referring to.
+It is obvious that, for every $k$ and $i\leq j$,
+(2.1) $\ell_{k+1,i}\geq\ell_{k,i},\hskip 28.45274pt\mbox{ and }\hskip 28.45274pt\ell_{k,j}\geq\ell_{k,i}.$
 We observe that describing a $2$-coloring is equivalent to describing a subset.
-Therefore, $\ell_{1,2}$, $\ell_{2,2}$, $\ell_{3,2}$ and $\ell_{6,2}$ coincide respectively with the numbers of regular, binary, ternary and good orbits on the power set.
-Restated in terms of colorings, Lemma 2.4 says that $\ell_{1,4}\geq 1$ for every primitive solvable group (one color is needed for the complement).
-A portion of the proof of Theorem 1.1 relies on showing that $\ell_{i,6-i}\geq 5$ for all primitive solvable groups and each $i=1,2,3$.
-The following combinatorial lemma is useful to estimate $\ell_{i,k}$ for large $k$.
+Therefore, $\ell_{2,1}$, $\ell_{2,2}$, $\ell_{2,3}$ and $\ell_{2,6}$ coincide respectively with the numbers of regular, $2$-regular, $3$-regular and good orbits on the power set.
+Restated in terms of colorings, Lemma 2.4 says that $\ell_{4,1}\geq 1$ for every primitive solvable group (one color is needed for the complement).
+A portion of the proof of Theorem 1.1 relies on showing that $\ell_{6-i,i}\geq 5$ for all primitive solvable groups and each $i=1,2,3$.
+The following combinatorial lemma is useful to estimate $\ell_{k,i}$ for large $k$.
 Lemma 2.5.
-Fix $i\in\{1,2,3,6\}$. If $\ell_{i,2}\geq 1$, then $\ell_{i,k}\geq{k\choose 2}$. If $\ell_{i,j}>\ell_{i,j-1}$ for some $3\leq j\leq k$, then $\ell_{i,k}\geq{k\choose j}$.
+Fix $k\geq 3$ and $i\in\{1,2,3,6\}$. If $\ell_{2,i}\geq 1$, then $\ell_{k,i}\geq{k\choose 2}$. If $\ell_{j,i}>\ell_{j-1,i}$ for some $3\leq j\leq k$, then $\ell_{k,i}\geq{k\choose j}$.
 Proof.
-The inequality $\ell_{i,j}>\ell_{i,j-1}$ means that there exists a suitable coloring that involves precisely $j$ colors.
+The inequality $\ell_{j,i}>\ell_{j-1,i}$ means that there exists a suitable coloring that involves precisely $j$ colors.
 Now suppose we have such a $j$-coloring.
@@
 ∎
-The following result requires a slightly deeper investigation. In the notation $\ell_{i,k}$, we write $\ell_{i,k}(G)$ to specify the group we are referring to.
+Corollary 2.6.
+Let $A$ be a primitive solvable group. Then $\ell_{5,1}(A)\geq 5$.
+Proof.
+Fix $A$.
+We have $\ell_{4,1}(A)\geq 1$ by Lemma 2.4.
+This means that $\ell_{2,1}(A)\geq 1$, or the inequality $\ell_{j,1}(A)>\ell_{j-1,1}(A)$ occurs at some index $j=3,4$.
+Considering the worst case, we obtain
+$\ell_{5,1}(A)\geq\min_{j=2,3,4}{5\choose j}=5,$
+where we used Lemma 2.5 with $k=5$, $i=1$. ∎
+The next result concerns $3$-colorings.
 Lemma 2.7.
-Let $A$ be a primitive solvable group. Then $\ell_{2,3}(A)\geq 5$.
+Let $A$ be a primitive solvable group. If $A\neq 1$, then $\ell_{3,2}(A)\geq 5$.
 Proof.
-Let $\Omega$ be the base set. By Lemma 2.2 and (2.1), we can assume $|\Omega|\leq 9$. Since we are interested in a lower bound on the number of orbits, it is enough to deal with the groups
+Let $\Omega$ be the base set, so that $|\Omega|\geq 2$.
+By Lemma 2.2 and (2.1), we can assume $|\Omega|\leq 9$.
+Since we are interested in a lower bound on the number of orbits, it is enough to deal with the solvable (over)groups
 […]
-It is easy to see that $\ell_{2,3}$ is at least $5$ in all these cases. In particular, using [GAP4], the command
+It is easy to see that $\ell_{3,2}$ is at least $5$ (actually $6$) in all these cases. In particular, using [GAP4], the command
 OrbitsDomain( A , Tuples([1..k], NrMovedPoints(A)) , Permuted )
 returns the orbits of $A$ on the $k$-colorings of $\Omega$. ∎
-We prove Theorem 1.1 through the following stronger statement, that lends itself to an inductive argument.
+2.3. Imprimitive groups
+We now explain the general strategy to deal with colorings of imprimitive groups.
+A similar idea is contained in [Ser96, Proof of Theorem 1.2], but we will go a little further.
+Let $A$ and $B$ be permutation groups on the finite sets $\Omega_{1}$ and $\Omega_{2}$ respectively.
+Let $G\leqslant B\wr_{|\Omega_{1}|}A$ act imprimitively on $\Omega=\Omega_{1}\times\Omega_{2}$.
+Let $x,y\geq 2$ be integers.
+Let $X=(X_{i})_{i=1}^{x}$ be a $x$-coloring of $\Omega_{1}$, namely $\Omega_{1}=\sqcup_{i=1}^{x}X_{i}$.
+Moreover, suppose we have $x$ inequivalent $y$-colorings of $\Omega_{2}$, say $Y_{1},\ldots,Y_{x}$.
+In particular, for each $i=1,\ldots,x$, we have that $\Omega_{2}=\sqcup_{j=1}^{y}Y_{i,j}$.
+The idea is to use the $y$-colorings of $\Omega_{2}$ as colors for $\Omega_{1}$.
+(In fact, we intentionally used the same index $i$ to count the colors in $X$ and the colorings of $\Omega_{2}$).
+We will construct a new $y$-coloring of $\Omega$, say $Z=(Z_{j})_{j=1}^{y}$, as follows.
+For each $j=1,\ldots y$, define
+$Z_{j}\>=\>\{(\omega_{1},\omega_{2})\in\Omega:\omega_{1}\in X_{i}\mbox{ and }\omega_{2}\in Y_{i,j}\mbox{ for some fixed }i\}.$
+It is easy to see that $(Z_{j})_{j=1}^{y}$ is actually a partition of $\Omega$.
+Lemma 2.8.
+The stabilizer of the coloring $Z$ in $G$ can be embedded in the semidirect product between the direct product of the stabilizers of the colorings $Y_{i}$’s in $B$ and the stabilizer of the coloring $X$ in $A$.
+Proof.
+First suppose $G=B\wr_{|\Omega_{1}|}A$. Since the colorings of $\Omega_{2}$ are inequivalent, the stabilizer of $Z$ is precisely the above semidirect product. The general case follows. ∎
+Lemma 2.9.
+Inequivalent $x$-colorings of $\Omega_{1}$ provide inequivalent $y$-colorings of $\Omega$.
+Proof.
+Fix $Y_{1},\ldots,Y_{x}$ colorings of $\Omega_{2}$, and let $X,X^{\prime}$ be two colorings of $\Omega_{1}$.
+Let $Z$ and $Z^{\prime}$ be the colorings of $\Omega=\Omega_{1}\times\Omega_{2}$ obtained from $X$ and $X^{\prime}$ respectively.
+If $Z$ and $Z^{\prime}$ lies in the same orbit of $B\wr_{|\Omega_{1}|}A$, then the induced colorings of $Z$ and $Z^{\prime}$ on $\Omega_{1}$ need to be in the same orbit of $A$.
+But these restrictions are precisely $X$ and $X^{\prime}$, and the proof follows.
+∎
+We are ready to prove Theorem 1.1. We actually prove the following stronger statement, that lends itself to an inductive argument:
 Theorem 2.10.
@@
 Primitive groups can be handled with Lemma 2.2 and Table 1.
-Let $\Omega=\Omega_{1}\times\Omega_{2}$, with $|\Omega_{1}|=a\geq 2$.
-Let $G\leqslant B\wr_{a}A$, with $A$ primitive on $\Omega_{1}$, and $B$ transitive on $\Omega_{2}$.
+Let $\Omega=\Omega_{1}\times\Omega_{2}$, with $|\Omega_{1}|\geq 2$.
+Let $G\leqslant B\wr_{|\Omega_{1}|}A$, with $A$ primitive on $\Omega_{1}$, and $B$ transitive on $\Omega_{2}$.
 By the inductive hypothesis applied to $\Omega_{2}$, we are in one of the following cases:
   • $B$ has at least $5$ good orbits on $\mathcal{P}(\Omega_{2})$.
-    We are going to use the good orbits of $B$ on $\mathcal{P}(\Omega_{2})$, say $\Sigma_{1}^{B},\ldots,\Sigma_{5}^{B}$, as colors for $\Omega_{1}$.
-Indeed, for each regular $5$-coloring of $\Omega_{1}$, say $\Omega_{1}=\sqcup_{j=1}^{5}\Delta_{j}$, we can construct a good orbit $\Delta^{G}$ for $G$ as follows:
-    $\Delta=\{(x,y)\in\Omega_{1}\times\Omega_{2}:x\in\Delta_{j}\mbox{ and }y\in\Sigma_{j}\mbox{ for some fixed }j\}.$
-    The setwise stabilizer of $\Delta$ in $G$ can be embedded in the semidirect product between the direct product of the stabilizers of the $\Sigma_{j}$’s in $B$ (which have orbits of length at most $6$), and the stabilizer of the regular $5$-coloring in $A$ (which is trivial).
-Inequivalent $5$-colorings of $\Omega_{1}$ provide inequivalent good orbits of $\Omega$.
-In particular, there are at least $\ell_{1,5}(A)$ good orbits of $G$.
-Now $\ell_{1,4}(A)\geq 1$ by Lemma 2.4, and so, from Lemma 2.5,
-    $\ell_{1,5}(A)\geq\min_{j=2,3,4}{5\choose j}=5.$
+    Since the goal is good $2$-colorings, we will use the construction above with $y=2$.
+In this case, set $x=5$.
+Let $\Delta_{1},\ldots,\Delta_{5}\subseteq\Omega_{2}$ be such that $\Delta_{1}^{B},\ldots,\Delta_{5}^{B}$ are distinct good orbits of $B$ on $\mathcal{P}(\Omega_{2})$.
+For each asymmetric $5$-coloring of $\Omega_{1}$, say $X=(X_{i})_{i=1}^{5}$, we construct a good orbit $\Sigma^{G}$ for $G$ as follows:
+    $\Sigma\>=\>\{(\omega_{1},\omega_{2})\in\Omega_{1}\times\Omega_{2}:\omega_{1}\in X_{i}\mbox{ and }\omega_{2}\in\Delta_{i}\mbox{ for some fixed }i\}.$
+    In fact, by Lemma 2.8, the set-stabilizer of $\Sigma$ in $G$ can be embedded in the semidirect product between the direct product of the set-stabilizers of the $\Delta_{i}$’s in $B$ (which have orbits of length at most $6$), and the stabilizer of the asymmetric $5$-coloring in $A$ (which is trivial).
+It follows from Lemma 2.9 that there are at least $\ell_{5,1}(A)$ good orbits of $G$.
+The proof of this part is completed by Corollary 2.6.
   • $B\cong\mathop{\mathrm{Sym}}(2)$.
     Here $\ell_{2,2}(B)=3$.
-Using the binary orbits of $B$, and the ternary $3$-colorings of $\Omega_{1}$, as before we can construct $\ell_{3,3}(A)$ good orbits of $G$.
-In fact, the length of any orbit will be bounded by $2\cdot 3=6$.
-Now $\ell_{3,3}(A)\geq\ell_{2,3}(A)\geq 5$ by (2.1) and Lemma 2.6.
+Using the $2$-asymmetric orbits of $B$, and the $3$-asymmetric $3$-colorings of $\Omega_{1}$, arguing as before we can construct $\ell_{3,3}(A)$ good orbits of $G$ (set $x=3$, $y=2$ in the construction above).
+In fact, note that orbit lengths are multiplied in the imprimitive action, so the length of each orbit will be bounded by $2\cdot 3=6$.
+Now $\ell_{3,3}(A)\geq\ell_{3,2}(A)\geq 5$ by (2.1) and Lemma 2.7.
   • $B\leqslant\mathop{\mathrm{Sym}}(3)$.
-    We can assume $B\cong\mathop{\mathrm{Sym}}(3)$, so that $\ell_{3,2}(B)=4$.
-Using the ternary orbits of $B$, and the binary $4$-colorings of $\Omega_{1}$, as before we can construct $\ell_{2,4}(A)$ good orbits of $G$.
-In fact, the length of any orbit will be bounded by $3\cdot 2=6$.
-Now $\ell_{2,4}(A)\geq\ell_{2,3}(A)\geq 5$ by (2.1) and Lemma 2.6.
-∎
-The cardinality of $\Delta$ in Theorem 1.1 grows linearly with the cardinality of $\Omega$, which is necessary as shown by $B\wr K$ when $B$ has some orbit of length exceeding $6$.
+    We can assume $B\cong\mathop{\mathrm{Sym}}(3)$, so that $\ell_{2,3}(B)=4$.
+Using the $3$-asymmetric orbits of $B$, and the $2$-asymmetric $4$-colorings of $\Omega_{1}$, as before we can construct $\ell_{4,2}(A)$ good orbits of $G$ (set $x=4$, $y=2$ in the construction above).
+The length of each orbit will be bounded by $3\cdot 2=6$.
+Finally, $\ell_{4,2}(A)\geq\ell_{3,2}(A)\geq 5$ by (2.1) and Lemma 2.7.
+The proof is complete. ∎
+The cardinality of the subset in Theorem 1.1 grows linearly with the cardinality of $\Omega$, which is necessary as shown by $B\wr K$, when $B$ has some orbit of length exceeding $6$, and the degree of $K$ grows.
 The bound on the derived length follows easily:
@@
 ∎
-The next example provides a transitive solvable group where no subset gives a metabelian stabilizer.
+In a previous version of this paper, Corollary 1.2 was obtained directly, without first bounding the length of the orbits.
+Indeed, the idea of proving a stronger statement by induction and gluing colorings is flexible enough to give structural results directly (see Theorem 2.16 below, for example).
+The next example provides a transitive solvable group where no subset gives a metabelian stabilizer:
 Example 2.11.
 Let $B=\mathop{\mathrm{Sym}}(4)\wr_{4}\mathop{\mathrm{Sym}}(4)$ act imprimitively on $\Omega=\{a,b,c,d\}\times\{1,2,3,4\}$.
-It is not hard to see that $B$ has a unique orbit on the power set whose setwise stabilizer is metabelian, namely $\Delta^{B}$, where
+It is not hard to see that $B$ has a unique orbit on the power set whose set-stabilizer is metabelian, namely $\Delta^{B}$, where
 $\Delta\>=\>\{(a,1),(b,1),(b,2),(b,3),(c,1),(c,2),(d,1),(d,2)\}.$
-The setwise stabilizer $H$ of $\Delta$ in $B$ is isomorphic to $\mathop{\mathrm{Sym}}(3)\times\mathop{\mathrm{Sym}}(3)\times D_{8}$.
+The set-stabilizer $H$ of $\Delta$ in $B$ is isomorphic to $\mathop{\mathrm{Sym}}(3)\times\mathop{\mathrm{Sym}}(3)\times(C_{2}\times C_{2})\wr_{2}C_{2}$.
 Now let $G=B\wr_{2}\mathop{\mathrm{Sym}}(2)$ act on $\Omega\times\{\alpha,\beta\}$.
-If we color both $\Omega\times\{\alpha\}$ and $\Omega\times\{\beta\}$ using $\Delta$, then the resulting setwise stabilizer will be isomorphic to $H\wr_{2}\mathop{\mathrm{Sym}}(2)$, whose derived length is $3$.
-Otherwise, we are forced to color a subset, say of $\Omega\times\{\alpha\}$, which does not provide a metabelian setwise stabilizer in $B$.
-So there is no hope of getting a metabelian setwise stabilizer in $G$.
-The next similar example involves $\mathrm{A\Gamma L}_{1}(8)$, and provides a transitive solvable group where no subset gives a stabilizer with all orbits of length at most $5$.
+If we color both $\Omega\times\{\alpha\}$ and $\Omega\times\{\beta\}$ using $\Delta$, then the resulting set-stabilizer will be isomorphic to $H\wr_{2}\mathop{\mathrm{Sym}}(2)$, whose derived length is $3$.
+Otherwise, we are forced to color a subset, say of $\Omega\times\{\alpha\}$, which does not provide a metabelian set-stabilizer in $B$.
+So there is no hope of getting a metabelian set-stabilizer in $G$.
+Since $\mathop{\mathrm{Sym}}(3)\wr_{2}\mathop{\mathrm{Sym}}(2)$ is not supersolvable, neither is $H\wr_{2}\mathop{\mathrm{Sym}}(2)$, and the same example shows that supersolvable set-stabilizers need not to exist.
+This is an improvement on [Glu24, Example 3.4], which deals with nilpotent set-stabilizers.
+The next similar construction involves $\mathrm{A\Gamma L}_{1}(8)$, and provides a transitive solvable group where no subset gives a stabilizer with all orbits of length at most $5$.
+We observe that, if a finite group acts on a set (even if not transitively), then the length of each orbit has to divide the order of the group.
+So, for example, $\mathrm{A\Gamma L}_{1}(8)$ does not have any orbit of length $5$.
 Example 2.12.
-We first notice that $\mathrm{A\Gamma L}_{1}(8)$ has precisely $4$ orbits on the power set whose stabilizers have orbits of length at most $4$.
+We first notice that $\mathrm{A\Gamma L}_{1}(8)$ has precisely $4$ orbits on the power set whose stabilizers have orbits of length at most $4$, and $4$ is attained by one of these orbits (a crucial fact is that $\ell_{2,3}(\mathrm{A\Gamma L}_{1}(8))=3$ in Table 1).
 Let $B=\mathrm{A\Gamma L}_{1}(8)\wr_{4}\mathop{\mathrm{Sym}}(4)$ act imprimitively on a set $\Omega$ of cardinality $32$.
-It is not hard to see that $B$ has a unique orbit $\Delta^{B}$ on the power set whose setwise stabilizer has orbits of length at most $5$ (actually $4$).
+It is not hard to see that $B$ has a unique orbit $\Delta^{B}$ on the power set where the set-stabilizer has orbits of length at most $5$ (actually $4$).
 Now let $G=B\wr_{2}\mathop{\mathrm{Sym}}(2)$ act on $\Omega\times\{\alpha,\beta\}$.
-If we color both $\Omega\times\{\alpha\}$ and $\Omega\times\{\beta\}$ using $\Delta$, then the resulting setwise stabilizer will have an orbit of length $4\cdot 2=8$.
-Otherwise, we are forced to color a subset, say of $\Omega\times\{\alpha\}$, which does not provide a setwise stabilizer (in $B$) with all orbits of length at most $5$.
+If we color both $\Omega\times\{\alpha\}$ and $\Omega\times\{\beta\}$ using $\Delta$, then the resulting set-stabilizer will have an orbit of length $4\cdot 2=8$.
+Otherwise, we are forced to color a subset, say of $\Omega\times\{\alpha\}$, which does not provide a set-stabilizer (in $B$) with all orbits of length at most $5$.
 So there is no hope of getting such a stabilizer in $G$.
 2.4. A result about $3$-colorings
-With the same ideas used in the proof of Theorem 2.7, it is possible to obtain the following result concerning $3$-colorings.
-We observe that Corollary 2.11 answers another question of Babai [Bab22, Question 7c].
+With the same ideas developed in the proof of Theorem 2.10, we obtain the following powerful result concerning $3$-colorings.
 Theorem 2.13.
-Let $G$ be a solvable permutation group on the finite set $\Omega$. Then there exist at least $5$ orbits of $3$-colorings of $\Omega$ whose stabilizers have all orbits of length at most $2$.
+Let $G$ be a solvable permutation group on the finite set $\Omega$, $|\Omega|\geq 2$.
+Then there exist at least $5$ orbits of $3$-colorings of $\Omega$ whose stabilizers have all orbits of length at most $2$.
 Proof.
-We have to prove that $\ell_{2,3}(G)\geq 5$.
-We can suppose that $G$ is transitive.
-Primitive groups can be handled with Lemma 2.6.
-Let $\Omega=\Omega_{1}\times\Omega_{2}$, with $|\Omega_{1}|=a\geq 2$.
-Let $G\leqslant B\wr_{a}A$, with $A$ primitive on $\Omega_{1}$, and $B$ transitive on $\Omega_{2}$.
-By the inductive hypothesis, we have that $\ell_{2,3}(B)\geq 5$.
-Arguing as in the proof of Theorem 2.7, using these binary $3$-colorings of $\Omega_{2}$ as colors for $\Omega_{1}$, and the regular $5$-colorings of $\Omega_{1}$, we can construct $\ell_{1,5}(A)$ binary $3$-colorings of $\Omega$.
-Finally, from Lemmas 2.4 and 2.5, we have $\ell_{1,5}(A)\geq 5$.
+We have to prove that $\ell_{3,2}(G)\geq 5$.
+We work by induction on $|\Omega|$, and we can suppose that $G$ is transitive.
+Also, primitive groups can be handled with Lemma 2.7.
+Let $\Omega=\Omega_{1}\times\Omega_{2}$, with $|\Omega_{1}|\geq 2$.
+Let $G\leqslant B\wr_{|\Omega_{1}|}A$, with $A$ primitive on $\Omega_{1}$, and $B$ transitive on $\Omega_{2}$.
+By the inductive hypothesis, we have that $\ell_{3,2}(B)\geq 5$.
+We argue as in Subsection 2.3, with $x=5$ and $y=3$.
+Using the $2$-asymmetric $3$-colorings of $\Omega_{2}$ as colors for $\Omega_{1}$, and the asymmetric $5$-colorings of $\Omega_{1}$, we can construct $\ell_{5,1}(A)$ inequivalent $2$-asymmetric $3$-colorings of $\Omega$.
+We have $\ell_{5,1}(A)\geq 5$ from Corollary 2.6, which makes the induction smooth and completes the proof.
 ∎
+The correspondent structural result answers [Bab22, Question 7c].
 Corollary 2.14.
 Let $G$ be a solvable permutation group on the finite set $\Omega$. Then there exists a $3$-coloring of $\Omega$ whose stabilizer is an elementary abelian $2$-group.
-As already noticed by Babai, not even a generic $2$-group can be obtained with just two colors: take $\mathop{\mathrm{Sym}}(4)\wr K$ for any solvable $K$ that is not a $2$-group.
-2.4. A CFSG-free proof for arbitrary primitive groups
+See [Glu24] for another positive result (with two colors) that emphatizes the importance of the prime $2$.
+In general, five colors are necessary and sufficient to achieve a trivial stabilizer [Ser96, Theorem 1.2].
+2.5. CFSG-free results for nonsolvable groups
 The foundation stone to prove a result like Theorem 1.1 is the fact that a primitive solvable group of large degree has a regular orbit on the power set [Glu83].
-With the trivial exceptions of the alternating and symmetric groups, the same is actually true in the nonsolvable case, but the known proofs of Cameron, Neumann, Saxl and Seress [CNS84, Ser97] require the classification of the finite simple groups.
-We observe that a more recent elementary result of Sun and Wilmes [SW15, SW16] has the following consequence.
+With the trivial exceptions of the alternating and symmetric groups, the same is actually true in the nonsolvable case, but the known proofs of Cameron, Neumann, Saxl and Seress [CNS84, Ser97] require the Classification of Finite Simple Groups.
+Actually, Cameron [Cam86] proved a stronger statement concerning the number of regular orbits.
+We observe that a more recent elementary result of Sun and Wilmes [SW15, SW16] is strong enough to provide an easier proof of Cameron’s theorem.
+[^1] Babai pointed out that a slightly weaker observation, which gives an elementary proof of [CNS84], was made by himself in [Bab22, Theorem 15.1].
 Theorem 2.15.
-Let $G$ be a primitive group not containing $\mathop{\mathrm{Alt}}(n)$. With only finitely many exceptions, $G$ has a regular orbit on the power set.
+Let $G$ be a primitive group of degree $n$ not containing $\mathop{\mathrm{Alt}}(n)$. Then the probability that a random orbit on the power set is a regular orbit tends to $1$ as $n$ grows.
 Proof.
-Let $n$ be the degree of $G$.
-We mimic the short proof in [CNS84], which uses the notions of base size and minimal degree in a clever way.
-At the crucial point, they obtain $|G|\geq 2^{\sqrt{n/2}}$, and use the classification to remain only with Johnson and Hamming groups, which are handled separetely.
-While Babai’s theorem [Bab81] is not enough, [SW16, Corollary 1.6] is sufficient for this job, having a thresold of $2^{n^{1/3}(\log n)^{7/3}}$.
+With the notation in [Cam86], it is sufficient to prove that $\varepsilon_{1}(G)\to 0$ when $n$ grows.
+We mimic Cameron’s proof, which uses the notions of base size and minimal degree in a clever way.
+At the crucial point, he obtains the desired result provided that $|G|\leq c^{\sqrt{n}}$, where $c$ is a constant (see the Remark at the end of page 308).
+Then, the Classification of Finite Simple groups is used to bound the order of a primitive group, remaining only with Johnson and Hamming groups, which are handled separately.
+While Babai’s theorem [Bab81] is not enough, [SW16, Corollary 1.6] is sufficient for this job, having a thresold of $d^{\,n^{1/3}(\log n)^{7/3}}$ where $d$ is a constant.
 ∎
-Let $G$ be an imprimitive group.
-To glue together the various pieces in Theorem 2.12, one has to be careful that large alternating groups are not induced on any subset.
-This fact makes the possible statements more complicated.
-One consequence is the following:
+We stress that Babai’s theorem [Bab81] has several applications, and Theorem 2.15 is a nice example where Sun-Wilmes improvement turns out to be crucial.
+(See [Ebe24] for another recent example.) We are now ready to prove Theorem 1.3, which answers [Bab22, Question 6a].
+Again, we prove a slightly stronger statement.
 Theorem 2.16.
-Let $G$ be a finite permutation group on $\Omega$ which does not involve $\mathop{\mathrm{Alt}}(n)$ for any $n\geq n_{0}$.
-Then there is a subset of $\Omega$ whose setwise stabilizer has all orbits of length bounded by a function of $n_{0}$.
+There exists a constant $C$ such that the following holds.
+Let $G$ be a permutation group on the finite set $\Omega$, $|\Omega|\geq 2$.
+Then there are at least two orbits of subsets of $\Omega$ whose setwise stabilizers have all nonalternating composition factors of order at most $C$.
+Proof.
+After Theorem 2.15, it is easy to see that such a constant $C$ exists for all primitive groups.
+Given $C$, we call good the orbits on the power set with the property in the statement.
+We can suppose that $G$ is transitive.
+By the Jordan-Hölder theorem, the property of having all nonalternating composition factors of order at most $C$, is a group property preserved by group extensions.
+Let $\Omega=\Omega_{1}\times\Omega_{2}$, with $|\Omega_{1}|\geq 2$.
+Let $G\leqslant B\wr_{|\Omega_{1}|}A$, with $A$ primitive on $\Omega_{1}$, and $B$ transitive on $\Omega_{2}$.
+By induction, $B$ admits two good orbits on $\mathcal{P}(\Omega_{2})$ (equivalently, two good $2$-colorings).
+Arguing as in Subsection 2.3 (set $x=y=2$), using these good orbits of $B$ as colors for $\Omega_{1}$, and the two good $2$-colorings of $\Omega_{1}$, we can construct two inequivalent good orbits of $G$ on $\mathcal{P}(\Omega)$.
+The proof is complete.
+∎
+With more work, and using the Classification of Finite Simple Groups, it should be possible to find an explicit version of Theorem 2.16 (it is enough to find $C$ for all primitive groups).
 3. Primitive groups with solvable stabilizer
+In this section, we move to consider pointwise stabilizers of very small sets.
 We would like to remark that, the idea of fixing a bounded number of points to obtain something interesting, is completely hopeless in imprimitive groups.
-This is a key difference with respect to setwise stabilizers (see Corollaries 1.2 and 2.11, for example).
+This is a key difference with respect to stabilizing sets (see Corollaries 1.2 or 2.14, for example).
 Example 3.1.
-Let $B$ be a transitive group with stabilizer $B_{0}$, and let $C_{k}$ be the cyclic group on $k$ points.
-Let $G=B\wr_{k}C_{k}$ act imprimitively on $|B:B_{0}|k$ points.
-The stabilizer of a point in $G$ is $H\cong B_{0}\times B^{k-1}$.
-For each $b\geq 1$, the intersection of any $b$ conjugates of $H$ will contain a copy of $B^{k-b}$, showing that fixing less than $k$ points is useless.
-Ultimately, the goal of this section is to prove Theorem 1.4.
-In the case of primitive solvable groups, this is [LS24, Theorem 1.1].
-The proof of this part is really about finding a good centralizer in every irreducible solvable linear group, which was done by Yang [Yan09] using ideas in [MW04].
-In general, as explained by Li and Zhang in [LZ11, Theorem 1.1], a primitive nonsolvable group with solvable stabilizer is almost simple or of a specific product type.
+Let $B$ be a transitive group with stabilizer $B_{0}$, and let $S\leqslant\mathop{\mathrm{Sym}}(d)$.
+Let $G=B\wr_{d}S$ act imprimitively on $|B:B_{0}|d$ points.
+The stabilizer of a point in $G$, say $H$, contains a copy of $B_{0}\times B^{d-1}$.
+For each $b\geq 1$, the intersection of any $b$ conjugates of $H$ will contain a copy of $B^{d-b}$, showing that fixing less than $d$ points is useless.
+Therefore, we will focus on primitive groups, with the goal of proving Theorem 1.5.
+This proof is more algebraic, but the combinatorial ideas in the previous section will also play a key role.
+In the case of primitive solvable groups, Theorem 1.5 is [LS24, Theorem 1.1].
+This part is really about finding a good centralizer in every irreducible solvable linear group, which was done by Yang [Yan09] using ideas in [MW04].
+In general, as explained by Li and Zhang in [LZ11, Sec.
+2], it is not hard to see that a primitive nonsolvable group with solvable stabilizer is almost simple or of a specific product type.
 3.1. Almost simple groups
@@
 10], together with a few properties of finite simple groups and linear groups.
-We report a strong version of the Schreier conjecture, which says that the group $\mathop{\mathrm{Out}}(T)$ of the outer automorphisms of a finite simple group is solvable.
+We first report a strong version of the Schreier conjecture, which says that the group $\mathop{\mathrm{Out}}(T)$ of the outer automorphisms of a finite simple group is solvable.
 Lemma 3.2.
@@
 Proof.
-This follows from the classification of the finite simple groups.
+This follows from the Classification of Finite Simple Groups.
 In particular, see [CCN+85, Pag.
@@
 Proof of Theorem 1.4.
-Let $T$ be the socle of $G$, and let $G_{0}$ be the normal subgroup containing $T$ which is described in [LZ11, Pag. 4]. We have that $H$ is an extension of $H_{0}=H\cap G_{0}$ with
+Let $T$ be the socle of $G$, and let $G_{0}$ be the normal subgroup containing $T$ which is described in [LZ11, Pag.
+4].
+We have that the maximal subgroup $H$ is an extension of $H_{0}=H\cap G_{0}$ with
 $H/H_{0}\cong G_{0}H/G_{0}\cong G/G_{0},$
@@
 To deal with (iii) uniformly, we use the fact that a solvable linear group has derived length bounded by a function on the rank [Dix68].
-The (relatively high) constant $10$ is attained by the Fischer group $Fi_{{}_{23}}$, which makes smooth the remaining computations.
+This and Lemma 3.2 prove Theorem 1.4 for some absolute constant.
+The (relatively high) constant $10$ is attained by the Fischer group $Fi_{{}_{23}}$, which makes smooth the remaining computations in (i)-(iii).
+A more careful inspection of the tables shows that $dl(H_{0})\leq 6$ in the infinite families of (iii), while we use a computer for the remaining single groups.
 We point out some mistakes in [LZ11, Table 15], and rely instead on the [GAP4] library CTblLib (v1.3.9).
-[^2] One mistake is in correspondence of the solvable maximal subgroups of $Fi_{{}_{23}}$.
-For example, a potential obstruction is given by a certain conjugacy class of solvable maximal subgroups $H$ of the baby monster, which seems to be not available on a computer.
+[^2] One mistake is in correspondence of $Fi_{{}_{23}}$.
+In fact, $Fi_{{}_{23}}$ has a unique conjugacy class of solvable maximal subgroups, namely AtlasSubgroup(‘‘Fi23’’,7)), with structure $43^{1+8}.2^{1+6}.3^{1+2}.2S$ 4 (with the ATLAS notation) and derived length $10$.
+A potential obstruction is given by a certain conjugacy class of solvable maximal subgroups $H$ of the baby monster, which seems to be not available on a computer.
 With the ATLAS notation [CCN+85], $H$ is isomorphic to $3^{11}.(S4\times 2S4)$.
-A group of order $3^{11}$ has nilpotency class at most $10$ and derived length at most $4$, showing that $dl(H)\leq 8$ in this case.
+Any group of order $3^{11}$ has nilpotency class at most $10$ and derived length at most $4$, showing that $dl(H)\leq 8$ in this case.
 ∎
 A weaker version of Theorem 1.4 says that there exists a constant $C$ such that if $G$ is a finite group with a solvable maximal subgroup $H$, and the derived length of $H$ is more than $C$, then $G$ is not simple.
-It would be interesting to find a proof of this fact which does not rely on the classification of the finite simple groups.
+It would be interesting to find a proof of this curious fact which does not rely on the Classification of Finite Simple Groups.
 3.2. Groups of product type
@@
 By [LZ11, Theorem 1.1], the primitive group $G$ that we care about satisfies
-$T\wr_{d}S\>\leqslant\>G\>\leqslant\>X\wr_{d}S.$
-(See also [Bur21, Sec.
-8.2].) Let $V\cong X^{d}$ be the base subgroup of $X\wr_{d}S$, and let $H<G$ be the stabilizer of a point in the product action.
+$T\wr_{d}S\>\leqslant\>G\>\leqslant\>X\wr_{d}S,$
+and acts on $|X:Y|^{d}$ points. (See also [Bur21, Sec. 8.2].)
+Proof of Theorem 1.5.
+Let $V\cong X^{d}$ be the base subgroup of $X\wr_{d}S$, and let $H<G$ be the stabilizer of a point in the product action.
 Let $W\cong Y^{d}$ be the subgroup of $V$ such that $H=WS\cap G$.
 Stabilizing two points corresponds to intersecting $H$ with one of its conjugates.
-To prove Theorem 1.4, we will find $v\in G\cap V$ such that the derived length of
+To prove the theorem, we will find $v\in G\cap V$ such that the derived length of
 $I_{v}\>=\>WS\cap(WS)^{v}$
-is bounded. The proof will follow, because $H\cap H^{v}\leqslant I_{v}$, and the derived length of a subgroup is bounded by the derived length of the group.
-Proof of Theorem 1.4.
-Let $\pi\colon G\to S$ be the natural projection, and let $v\in V$. Since $Ker(\pi)=G\cap V$, we have
-$dl(I_{v})\leq dl(I_{v}\cap V)+dl(\pi(I_{v})).$
+is bounded by an absolute constant (actually $13$).
+The proof will follow, because $H\cap H^{v}\leqslant I_{v}$, and the derived length of a subgroup is bounded by the derived length of the group.
+Let $\pi\colon G\to S$ be the natural projection, and let $v\in V$.
+Since $Ker(\pi)=G\cap V$, we have
+(3.1) $dl(I_{v})\>\leq\>dl(I_{v}\cap V)+dl(\pi(I_{v})).$
 We observe that
@@
 $dl(I_{v}\cap V)\leq dl(W)=dl(Y).$
-Now $dl(Y)$ is bounded by Theorem 1.3, and the rest of the proof deals with $\pi(I_{v})$.
+Now $dl(Y)\leq 10$ by Theorem 1.4, and the rest of the proof deals with $\pi(I_{v})$.
 It is not hard to see that $C_{S}(v)=S\cap S^{v}\subseteq\pi(I_{v})$, but in general $\pi(I_{v})$ can be quite bigger.
@@
 $Y\cap x_{i}^{-1}Yx_{\sigma(i)}\neq\emptyset\hskip 28.45274pt\mbox{ if and onlyif }\hskip 28.45274ptx_{\sigma(i)}\in Yx_{i}Y,$
-we obtain
-(3.1) $\pi(I_{v})=\{\sigma\in S:\>x_{\sigma(i)}\in Yx_{i}Y\mbox{ for each }i=1,\ldots,d\}.$
-We are going to look at the double cosets of $Y$ as colors. Let $c_{1},\ldots,c_{r}\in X$ be representatives of these double cosets. For any $v\in V$, and $j\in\{1,\ldots,r\}$, let
+we obtain the key equality
+(3.2) $\pi(I_{(x_{1},\ldots,x_{d})})\>=\>\{\sigma\in S:\>x_{\sigma(i)}\in Yx_{i}Y\mbox{ for each }i=1,\ldots,d\}.$
+This suggests to look at the double cosets of $Y$ in $X$ as colors, and connects Theorem 1.5 with the matter of Section 2.
+Let $c_{1},\ldots,c_{r}\in X$ be representatives of these double cosets ($r\geq 2$).
+For any $v\in V$, and $j\in\{1,\ldots,r\}$, let
 $\Delta_{j}=\{i\in\{1,\ldots,d\}:x_{i}\in Yc_{j}Y\}.$
 Of course the $\Delta_{j}$’s form a partition of $\{1,\ldots,d\}$. The equality in (3.2) shows that $\pi(I_{v})$ only depends on this partition, and in particular
-(3.2) $\pi(I_{v})=\cap_{j=1}^{r}Stab_{S}(\Delta_{j})=\cap_{j=1}^{r-1}Stab_{S}(\Delta_{j}).$
+(3.3) $\pi(I_{v})\>=\>\cap_{j=1}^{r}Stab_{S}(\Delta_{j})\>=\>\cap_{j=1}^{r-1}Stab_{S}(\Delta_{j}).$
 To construct $v$, we will use only two elements $c_{1}$ and $c_{2}$ (and so two sets $\Delta_{1}$ and $\Delta_{2}=\{1,\ldots,d\}\setminus\Delta_{1}$).
@@
 This is assured if we choose representatives $c_{i}$’s which are in $T$.
-We can surely choose $c_{1}=1$ and $c_{2}\in T\setminus\{1\}$, because otherwise $T\leqslant Y$, which is impossible because $Y$ is solvable.
-Finally, choosing $\Delta_{1}=\Delta\subseteq\{1,\ldots,d\}$ as provided by Corollary 1.2, from (3.2) we get $dl(\pi(I_{v}))\leq 3$, concluding $dl(I_{v})\leq 13$ as desired.
+We can surely choose $c_{1}=1$ and $c_{2}\in T\setminus\{1\}$, because otherwise $T\leqslant Y$, which is impossible because $Y$ is core-free.
+Finally, if $\Delta_{1}=\Delta\subseteq\{1,\ldots,d\}$ is as provided by Corollary 1.2, from (3.3) we get $dl(\pi(I_{v}))\leq 3$, concluding by (3.1) that $dl(I_{v})\leq 13$ as desired.
 ∎
@@
 As explained, one reason is that we want $v\in G$, but there is also a different motivation.
-In fact, it is well known that the number $r$ of double cosets of $Y$ in $X$ equals the rank of the primitive group $X$ with solvable stabilizer $Y$, and there actually exist many doubly transitive groups of this type (as $\mathop{\mathrm{Sym}}(5)$, for example).
-A more detailed analysis of highly transitive almost simple groups, and the additional use of Corollary 2.11, might help to find the right constant in Theorem 1.4.
-We remark that this constant is not even known in the case of solvable groups, where it lies somewhere between $4$ and $9$ (see [LS24, Remark 2.9]).
+In fact, it is well known that the number $r$ of double cosets of $Y$ in $X$ equals the rank of the primitive group $X$ with solvable stabilizer $Y$, and there actually exist some doubly transitive groups of this type (as $\mathop{\mathrm{Sym}}(5)$, for example).
+A more detailed analysis of almost simple groups, and the additional use of Corollary 2.14, might help to find the right constant in Theorem 1.5.
+In the case of primitive solvable groups, this lies somewhere between $4$ and $9$ (see [LS24, Remark 2.9]).
+H.Y.
+Huang pointed out that $5$ is attained by the almost simple group $\mathop{\mathrm{Aut}}(\mathrm{P}\Omega^{+}_{8}(3))$.
+With some more work, it is not hard to use this information to construct a group of product type where the constant $6$ is achieved.
+A small corollary on set-stabilizers can be deduced from Theorem 1.5.
 Corollary 3.5.
@@
 If $\{x_{1},x_{2}\}\subseteq\Delta$, then $C_{G}(\Delta)$ is solvable of bounded derived length.
-If $|\Delta|\leq 4$, then $Stab_{G}(\Delta)/C_{G}(\Delta)\leqslant\mathop{\mathrm{Sym}}(4)$, and it follows that $Stab_{G}(\Delta)$ is solvable of bounded derived lenght.
+Now $Stab_{G}(\Delta)/C_{G}(\Delta)\leqslant\mathop{\mathrm{Sym}}(4)$, and it follows that $Stab_{G}(\Delta)$ is solvable of bounded derived lenght.
 ∎

### v4

@@
 We prove various results about partitions of $\Omega$ whose stabilizers have good properties.
-In particular, we obtain the following three theorems: In every solvable permutation group, there is a subset whose stabilizer has all orbits of length at most $6$, which is best possible and answers two questions of Babai.
+In particular, in every solvable permutation group there is a set-stabilizer whose orbits have length at most $6$, which is best possible and answers two questions of Babai.
 Every solvable maximal subgroup of any almost simple group has derived length at most $10$, which is best possible.
-In every primitive group with solvable stabilizer, there are two points whose pointwise stabilizer has derived length bounded by an absolute constant.
+In every primitive group with solvable stabilizer, there are two points whose stabilizer has derived length bounded by an absolute constant.
 Key words and phrases:
@@
 In this paper, we are interested in a small variation on these questions.
-More precisely, we fix only the bare minimum (so one subset, or a pair of points) and we do not aim to find a trivial stabilizer, since this would be impossible in general, but to find a stabilizer that has at least some good properties.
-We will see that this can be done in a few interesting situations.
-Regarding subsets, Babai [Bab22] has proved that in every solvable permutation group there exists a set-stabilizer whose orbits have length bounded by a nonexplicit absolute constant.
-Answering [Bab22, Question 7a], our first main result provides the best possible constant:
+More precisely, we fix only the bare minimum (so one subset, or a pair of points) and we do not aim to find a trivial stabilizer, as that would be impossible in general, but to find a stabilizer that has at least some good properties.
+Regarding subsets, Babai [Bab22] has showed that in every solvable permutation group there exists a set-stabilizer whose orbits have length bounded by a nonexplicit absolute constant.
+Answering [Bab22, Quest.
+7b], our first main result provides the best possible constant:
 Theorem 1.1.
@@
 Then there exists a subset of $\Omega$ whose setwise stabilizer has all orbits of length at most $6$.
-The constant $6$ is the best possible.
+The constant $6$ is best possible.
 The following consequence is immediate.
@@
 Let $G\leqslant\mathop{\mathrm{Sym}}(\Omega)$ be a solvable permutation group. Then there exists a subset of $\Omega$ whose setwise stabilizer has derived length at most $3$.
-It happens that metabelian stabilizers need not to exist in general (see Example 2.11), so Corollary 1.2 answers [Bab22, Question 7b].
-Incidentally, the same example answers negatively a question in [Glu24] on the existence of supersolvable set-stabilizers.
-We stress that the two constants $6$ and $3$ above turned out to be a little higher than what initially expected, and better results can be obtained for primitive solvable groups.
+It happens that metabelian stabilizers need not exist in general (see Example 2.11), so Corollary 1.2 answers [Bab22, Quest.
+7a].
+Incidentally, the same example answers negatively a question in [Glu25] on the existence of supersolvable set-stabilizers.
+We stress that the two constants $6$ and $3$ above turned out to be higher than initially expected, and better results can be obtained for primitive solvable groups.
 As far as we know, Theorem 1.1 is the first statement concerning stabilizers of imprimitive groups where the property involved is not preserved by a group extension, and part of the difficulty of the proof is in finding the right statement to prove by induction.
+This strategy, combined with a detailed analysis of primitive groups, turned out to be both powerful and flexible.
 With similar ideas, we are able to show that, in every solvable permutation group, stabilizing two disjoint subsets can force the orbits to have length at most $2$.
-This gives a strong positive answer to [Bab22, Question 7c].
-Moreover, we give some elementary proofs concerning nonsolvable groups: we provide an easier proof of a theorem of Cameron [Cam86], and answer [Bab22, Question 6a].
+This gives a strong positive answer to [Bab22, Quest.
+7c].
+Moreover, we give elementary proofs concerning nonsolvable groups: we provide an easier proof of a theorem of Cameron [Cam86], and answer [Bab22, Quest.
+6a].
 In particular, we prove the following:
 Theorem 1.3.
-Let $G$ be any permutation group on a finite set. Then there exists a setwise stabilizer whose nonalternating composition factors have order bounded by an absolute constant.
+Let $G$ be any permutation group on a finite set. Then there exists a set-stabilizer whose nonalternating composition factors have order bounded by an absolute constant.
 This is new even using the Classification of Finite Simple Groups.
 We come to the question of fixing two points.
-Extending the work of Seress [Ser96], Burness [Bur21] showed that in a primitive group with solvable stabilizer, there always exist $5$ points whose pointwise stabilizer is trivial.
-The hypothesis of having a solvable stabilizer is natural, if we aim to structural results on the stabilizer of two points.
-For primitive solvable groups, in [LS24] the authors observed that there exist two points whose pointwise stabilizer has bounded derived length.
+Extending the work of Seress [Ser96], Burness [Bur21] showed that in a primitive group with solvable stabilizer there always exist $5$ points whose pointwise stabilizer is trivial.
+The hypothesis of having a solvable stabilizer is natural, if we aim to find structural results on the stabilizer of two points.
+For primitive solvable groups, in [LS25] the authors observed that there exist two points whose pointwise stabilizer has bounded derived length.
 On the other hand, motivated by various applications, Li and Zhang [LZ11] have classified the solvable maximal subgroups of the almost simple groups.
@@
 If $H$ is solvable, then the derived length of $H$ is at most $10$.
-The constant $10$ is attained only by the solvable maximal subgroups of the Fischer group $Fi_{{}_{23}}$.
+The constant $10$ is attained only by the solvable maximal subgroups of the Fischer group $\mathop{\mathrm{Fi}}_{{}_{23}}$.
 Despite Theorem 1.4, it is easy to exhibit primitive nonsolvable groups of product type with solvable stabilizer of arbitrarily large derived length.
-Dealing with the product type case, we are able to answer in the affirmative the Question at the end of [LS24]:
+Dealing with the product type case, we are able to answer in the affirmative the question at the end of [LS25]:
 Theorem 1.5.
 Let $G$ be a primitive group with solvable stabilizer. Then there exist two points whose pointwise stabilizer has derived length bounded by an absolute constant.
-The proof of Theorem 1.5 is an application of Corollary 1.2, and it is perhaps surprising that, to answer a question about pointwise stabilizers of primitive nonsolvable groups, one has to deal with set-stabilizers of arbitrary solvable permutation groups.
-The idea, usually applied in the context of imprimitive linear groups, is using a partition to construct a vector with small centralizer (see [MW04, Sec.
-4]).
-Our argument involves a double coset decomposition and a neat bound gives the constant $13$.
-It is likely that a more in-depth study of almost simple groups leads to an improvement, but the best possible constant is not even known for solvable groups [LS24].
-The paper is organized as follows.
-In Section 2 we study set-stabilizers, proving Theorem 1.1 and the other related results.
+The proof of Theorem 1.5 is an application of Corollary 1.2 (or of the nonexplicit result of Babai), and it is perhaps surprising that, to answer a question about pointwise stabilizers of primitive nonsolvable groups, one has to deal with set-stabilizers of arbitrary solvable permutation groups.
+The idea, usually applied in the context of imprimitive linear groups, is using a partition to construct a vector with small centralizer (see the book [MW93] for various examples).
+Our argument involves a double coset decomposition, and a crude bound gives the constant $13$.
+It is likely that a more in-depth study of almost simple groups leads to an improvement, but the best possible constant is not even known for solvable groups [LS25, Rem.
+2.9].
+The paper is simply organized.
+In Section 2 we study set-stabilizers, proving Theorems 1.1 and 1.3 and other related results.
 In Section 3, we prove Theorems 1.4 and 1.5.
@@
 Skresanov.
-The first asked for a structural result in the spirit of Corollary 2.14 (see also [Glu24]).
-The second noticed that the (relatively high) constant $10$ in Theorem 1.4 is attained by $Fi_{{}_{23}}$, and that the constant in Theorem 1.5 is at least $5$ for almost simple groups.
+The first asked for a structural result in the spirit of Corollary 2.14 (see also [Glu25]).
+The second noticed that the constant $10$ in Theorem 1.4 is attained by $\mathop{\mathrm{Fi}}_{{}_{23}}$, and that the constant in Theorem 1.5 is at least $5$ for almost simple groups.
 The third brought the article [Bab22] to my attention, and helped me with the computer calculations in Lemma 2.7.
@@
 When $G$ acts on a finite set $\Omega$, then $G$ acts naturally on the power set $\mathcal{P}(\Omega)$.
-If $\Delta\subseteq\Omega$, we write $C_{G}(\Delta)$ for the pointwise stabilizer, and $Stab_{G}(\Delta)$ for the setwise stabilizer.
+If $\Delta\subseteq\Omega$, we write $C_{G}(\Delta)$ for the pointwise stabilizer, and $\mathop{\mathrm{Stab}}_{G}(\Delta)$ for the setwise stabilizer.
 For each $g\in G$, it is easy to see that the stabilizers of $\Delta$ and $\Delta^{g}$ are conjugate subgroups.
-As usual, we say that $\Delta$ lies in a regular orbit (i.e., the orbit $\Delta^{G}$) if $Stab_{G}(\Delta)=1$.
+As usual, we say that $\Delta$ lies in a regular orbit (i.e., the orbit $\Delta^{G}$) if $\mathop{\mathrm{Stab}}_{G}(\Delta)=1$.
 Of course, as subgroups of $G$, the stabilizers still act on $\Omega$.
 Definition.
-For $i\geq 1$, we say that $\Delta\subseteq\Omega$ lies in a $i$-regular orbit if all the orbits of $Stab_{G}(\Delta)$ on $\Omega$ have length at most $i$.
+For $i\geq 1$, we say that $\Delta\subseteq\Omega$ lies in a $i$-regular orbit if all the orbits of $\mathop{\mathrm{Stab}}_{G}(\Delta)$ on $\Omega$ have length at most $i$.
 The $1$-regular orbits are precisely the genuine regular orbits.
@@
 (This is motivated by Theorem 1.1.) We note that, by definition,
-[…]
+ $\mbox{ regular }\Rightarrow\mbox{ 2-regular }\Rightarrow\mbox{ 3-regular }\Rightarrow\mbox{ good }.$
 Remark 2.1.
@@
 We now make a simple consideration. If $G\leqslant\mathop{\mathrm{Sym}}(\Omega)$ and $\Delta=\{x_{1},\ldots,x_{b}\}$ is a base (i.e., $C_{G}(\Delta)=1$), then
-$Stab_{G}(\Delta)\cong Stab_{G}(\Delta)/C_{G}(\Delta)\leqslant\mathop{\mathrm{Sym}}(b).$
-So, if $Stab_{G}(\Delta)$ is solvable, then $dl(Stab_{G}(\Delta))\leq f(b)$, where
+$\mathop{\mathrm{Stab}_{G}}(\Delta)\cong\mathop{\mathrm{Stab}_{G}}(\Delta)/C_{G}(\Delta)\leqslant\mathop{\mathrm{Sym}}(b).$
+So, if $\mathop{\mathrm{Stab}}_{G}(\Delta)$ is solvable, then $dl(\mathop{\mathrm{Stab}}_{G}(\Delta))\leq f(b)$, where
 $f(b)=\>\max\,\{dl(S):S\leqslant\mathop{\mathrm{Sym}}(b),S\mbox{ solvable}\}.$
@@
 Proof.
-The first part follows from [MW04, Lemma 4.1(ii)]. The remaining cases have been handled with the library PrimitiveGroups in [GAP4]. ∎
-$\dfrac{1^{1}}{1_{1_{1}}}$ GAP Id Regular orbits on $\mathcal{P}(\Omega)$ $2$-regular orbits on $\mathcal{P}(\Omega)$ $3$-regular orbits on $\mathcal{P}(\Omega)$ Good orbits on $\mathcal{P}(\Omega)$
+The first part follows from [MW04, Lem. 4.1(ii)]. The remaining cases have been handled with the library PrimitiveGroups in [GAP4]. ∎
+$\dfrac{1^{1}}{1_{1_{1}}}$ GAP Id Regular orbits on $\mathcal{P}(\Omega)$ 22-regular orbits on $\mathcal{P}(\Omega)$ 33-regular orbits on $\mathcal{P}(\Omega)$ Good orbits on $\mathcal{P}(\Omega)$
 $\ell_{2,1}$ $\ell_{2,2}$ $\ell_{2,3}$ $\ell_{2,6}$
-$\dfrac{1^{1}}{1_{1_{1}}}$ $[2,1]$ $1$ $3$ $3$ $3$
-$\dfrac{1^{1}}{1_{1_{1}}}$ $[3,1]$ $2$ $2$ $4$ $4$
-$\dfrac{1^{1}}{1_{1_{1}}}$ $[3,2]$ 00 $2$ $4$ $4$
-$\dfrac{1^{1}}{1_{1_{1}}}$ $[4,1]$ 00 $1$ $3$ $5$
-$\dfrac{1^{1}}{1_{1_{1}}}$ $[4,2]$ 00 $1$ $3$ $5$
-$\dfrac{1^{1}}{1_{1_{1}}}$ $[5,2]$ 00 $6$ $6$ $8$
-$\dfrac{1^{1}}{1_{1_{1}}}$ $[5,3]$ 00 $2$ $2$ $6$
-$\dfrac{1^{1}}{1_{1_{1}}}$ $[7,2]$ $2$ $16$ $16$ $16$
-$\dfrac{1^{1}}{1_{1_{1}}}$ $[7,3]$ $4$ $4$ $10$ $10$
-$\dfrac{1^{1}}{1_{1_{1}}}$ $[7,4]$ 00 $4$ $6$ $8$
-$\dfrac{1^{1}}{1_{1_{1}}}$ $[8,1]$ $3$ $5$ $5$ $6$
-$\dfrac{1^{1}}{1_{1_{1}}}$ $[8,2]$ 00 00 $3$ $6$
-$\dfrac{1^{1}}{1_{1_{1}}}$ $[9,2]$ 00 $10$ $10$ $24$
-$\dfrac{1^{1}}{1_{1_{1}}}$ $[9,3]$ $4$ $6$ $6$ $14$
-$\dfrac{1^{1}}{1_{1_{1}}}$ $[9,4]$ $4$ $8$ $8$ $12$
-$\dfrac{1^{1}}{1_{1_{1}}}$ $[9,5]$ 00 $4$ $4$ $12$
-$\dfrac{1^{1}}{1_{1_{1}}}$ $[9,6]$ 00 00 $4$ $10$
-$\dfrac{1^{1}}{1_{1_{1}}}$ $[9,7]$ 00 00 $4$ $10$
+$\dfrac{1^{1}}{1_{1_{1}}}$ $[2,1]$ 11 33 33 33
+$\dfrac{1^{1}}{1_{1_{1}}}$ $[3,1]$ 22 22 44 44
+$\dfrac{1^{1}}{1_{1_{1}}}$ $[3,2]$ 0 22 44 44
+$\dfrac{1^{1}}{1_{1_{1}}}$ $[4,1]$ 0 11 33 55
+$\dfrac{1^{1}}{1_{1_{1}}}$ $[4,2]$ 0 11 33 55
+$\dfrac{1^{1}}{1_{1_{1}}}$ $[5,2]$ 0 66 66 88
+$\dfrac{1^{1}}{1_{1_{1}}}$ $[5,3]$ 0 22 22 66
+$\dfrac{1^{1}}{1_{1_{1}}}$ $[7,2]$ 22 $16$ $16$ $16$
+$\dfrac{1^{1}}{1_{1_{1}}}$ $[7,3]$ 44 44 $10$ $10$
+$\dfrac{1^{1}}{1_{1_{1}}}$ $[7,4]$ 0 44 66 88
+$\dfrac{1^{1}}{1_{1_{1}}}$ $[8,1]$ 33 55 55 66
+$\dfrac{1^{1}}{1_{1_{1}}}$ $[8,2]$ 0 0 33 66
+$\dfrac{1^{1}}{1_{1_{1}}}$ $[9,2]$ 0 $10$ $10$ $24$
+$\dfrac{1^{1}}{1_{1_{1}}}$ $[9,3]$ 44 66 66 $14$
+$\dfrac{1^{1}}{1_{1_{1}}}$ $[9,4]$ 44 88 88 $12$
+$\dfrac{1^{1}}{1_{1_{1}}}$ $[9,5]$ 0 44 44 $12$
+$\dfrac{1^{1}}{1_{1_{1}}}$ $[9,6]$ 0 0 44 $10$
+$\dfrac{1^{1}}{1_{1_{1}}}$ $[9,7]$ 0 0 44 $10$
 Table 1. Primitive solvable groups with less than $5$ regular orbits on $\mathcal{P}(\Omega)$.
@@
 Proof.
-As before, using [MW04, Lemma 4.1(ii)], and the library PrimitiveGroups in [GAP4].
-We observe that $\mathrm{A\Gamma L}_{1}(8)$ and PrimitiveGroup(9,6) do not have any $2$-regular orbit on $\mathcal{P}(\Omega)$, but still have some abelian set-stabilizer.
+As before, using [MW04, Lem.
+4.1(ii)], and the library PrimitiveGroups in [GAP4].
+We observe that $\mathrm{A\Gamma L}_{1}(8)$ and PrimitiveGroup(9,6) do not have any 22-regular orbit on $\mathcal{P}(\Omega)$, but still have some abelian set-stabilizer.
 ∎
@@
 If $G\leqslant\mathop{\mathrm{Sym}}(\Omega)$ is a primitive solvable group, then there exist disjoint subsets $\Delta_{1},\Delta_{2},\Delta_{3}$ such that
-[…]
+$\cap_{j=1}^{3}\,\mathop{\mathrm{Stab}_{G}}(\Delta_{j})\>=\>1.$
 Proof.
-This is [MW04, Lemma 4.1(iii)]. As before, the proof works by checking the finitely many cases with $|\Omega|\leq 9$. ∎
-Lemma 2.4 is sharp, as shown by $\mathop{\mathrm{Sym}}(4)$.
+This is [MW04, Lem. 4.1(iii)]. ∎
 2.2. Good colorings
-The standard argument to deal with imprimitive groups is [Glu83, Theorem 2], which was used by Gluck to show that permutation groups of odd order have a regular orbit on the power set.
-This does not work with a property such as “have all orbits of length at most $C$”, where $C$ is a constant.
+The standard argument to deal with imprimitive groups is [Glu83, Th.
+2], which was used by Gluck to show that permutation groups of odd order have a regular orbit on the power set.
+This does not work with a property such as “have all orbits of length at most CC”, where $C$ is a constant.
 We use a different method based on good orbits and partitions that we are going to explain.
-A key observation, already in the lines of the proof of [Bab22, Theorem 12.1], is that studying stabilizers of partitions can also be useful to obtain results concerning stabilizers of a single subset.
-We need some more definitions.
-An asymmetric coloring of a permutation group $G$ on $\Omega$ is a partition of $\Omega$ with the property that only the identity fixes every part.
+A key observation, already in the lines of the proof of [Bab22, Th.
+12.1], is that studying stabilizers of partitions can also be useful to obtain results concerning stabilizers of a single subset.
+Fix $G\leqslant\mathop{\mathrm{Sym}}(\Omega)$, and let $j]=\{1,\ldots,j\}$ (the set of colors).
+A coloring is a function $\Omega\to[j]$, and a $k$-coloring is a coloring with at most $k$ colors.
+(This can also be described by the preimages of the colors, which form a partition of $\Omega$.) A coloring is asymmetric if only the identity fixes every color.
 We refer the reader to [Bab22, Sec.
 2] for some history concerning breaking-symmetry results and some conflicting terminology.
-We refine this notion as follows:
 Definition.
-Let $G\leqslant\mathop{\mathrm{Sym}}(\Omega)$, and $k\geq 2$.
-A $k$-coloring is a coloring of $\Omega$ with at most $k$ colors.
 For $i\geq 1$, an $i$-asymmetric $k$-coloring is a $k$-coloring such that the stabilizer of the coloring in $G$ has all orbits on $\Omega$ of length at most $i$.
@@
 The total number of $k$-colorings of $\Omega$ is $k^{|\Omega|}$, but many of these are equivalent under the action of $G$.
-Fixed $G\leqslant\mathop{\mathrm{Sym}}(\Omega)$, for each $k\geq 2$ we write
-[…]
-for the numbers of $G$-orbits of asymmetric, $2$-asymmetric, $3$-asymmetric and good $k$-colorings.
+Fix $G\leqslant\mathop{\mathrm{Sym}}(\Omega)$; for each $k\geq 2$ we write
+$\ell_{k,1},\hskip 14.22636pt\ell_{k,2},\hskip 14.22636pt\ell_{k,3},\hskip 14.22636pt\ell_{k,6},$
+for the numbers of $G$-orbits of asymmetric, $2$-asymmetric, 33-asymmetric and good $k$-colorings.
 When necessary, we write $\ell_{k,i}(G)$ to stress the group $G$ we are referring to.
@@
 We observe that describing a $2$-coloring is equivalent to describing a subset.
-Therefore, $\ell_{2,1}$, $\ell_{2,2}$, $\ell_{2,3}$ and $\ell_{2,6}$ coincide respectively with the numbers of regular, $2$-regular, $3$-regular and good orbits on the power set.
+Therefore, $\ell_{2,1}$, $\ell_{2,2}$, $\ell_{2,3}$ and $\ell_{2,6}$ coincide respectively with the numbers of regular, $2$-regular, 33-regular and good orbits on the power set.
 Restated in terms of colorings, Lemma 2.4 says that $\ell_{4,1}\geq 1$ for every primitive solvable group (one color is needed for the complement).
@@
 Proof.
-The inequality $\ell_{j,i}>\ell_{j-1,i}$ means that there exists a suitable coloring that involves precisely $j$ colors.
+The strict inequality $\ell_{j,i}>\ell_{j-1,i}$ means that there exists a suitable coloring that involves precisely $j$ colors.
 Now suppose we have such a $j$-coloring.
@@
 Proof.
-Fix $A$.
+Fix AA.
 We have $\ell_{4,1}(A)\geq 1$ by Lemma 2.4.
@@
 Since we are interested in a lower bound on the number of orbits, it is enough to deal with the solvable (over)groups
-[…]
+$\mathop{\mathrm{Sym}}(2),\hskip 5.69046pt\mathop{\mathrm{Sym}}(3),\hskip 5.69046pt\mathop{\mathrm{Sym}}(4),\hskip 5.69046pt\mathrm{AGL}_{1}(5),\hskip 5.69046pt\mathrm{AGL}_{1}(7),\hskip 5.69046pt\mathrm{A\Gamma L}_{1}(8),\hskip 5.69046pt\mathrm{AGL}_{2}(3).$
 It is easy to see that $\ell_{3,2}$ is at least $5$ (actually $6$) in all these cases. In particular, using [GAP4], the command
@@
 We now explain the general strategy to deal with colorings of imprimitive groups.
-A similar idea is contained in [Ser96, Proof of Theorem 1.2], but we will go a little further.
 Let $A$ and $B$ be permutation groups on the finite sets $\Omega_{1}$ and $\Omega_{2}$ respectively.
@@
 In particular, for each $i=1,\ldots,x$, we have that $\Omega_{2}=\sqcup_{j=1}^{y}Y_{i,j}$.
-The idea is to use the $y$-colorings of $\Omega_{2}$ as colors for $\Omega_{1}$.
+The point is to use the $y$-colorings of $\Omega_{2}$ as colors for $\Omega_{1}$.
 (In fact, we intentionally used the same index $i$ to count the colors in $X$ and the colorings of $\Omega_{2}$).
@@
 $Z_{j}\>=\>\{(\omega_{1},\omega_{2})\in\Omega:\omega_{1}\in X_{i}\mbox{ and }\omega_{2}\in Y_{i,j}\mbox{ for some fixed }i\}.$
-It is easy to see that $(Z_{j})_{j=1}^{y}$ is actually a partition of $\Omega$.
+It is easy to see that $(Z_{j})_{j=1}^{y}$ is a partition of $\Omega$.
 Lemma 2.8.
-The stabilizer of the coloring $Z$ in $G$ can be embedded in the semidirect product between the direct product of the stabilizers of the colorings $Y_{i}$’s in $B$ and the stabilizer of the coloring $X$ in $A$.
+The stabilizer of the coloring $Z$ in $G$ can be embedded in the semidirect product between the direct product of the stabilizers of the colorings $Y_{i}$ in $B$ and the stabilizer of the coloring $X$ in $A$.
+More precisely,
+$\cap_{j}\mathop{\mathrm{Stab}_{G}}(Z_{j})\>\leqslant\>\left(\prod_{i}\left(\cap_{j}\mathop{\mathrm{Stab}_{B}}(Y_{i,j})\right)\right)\rtimes\left(\cap_{i}\mathop{\mathrm{Stab}_{A}}(X_{i})\right).$
 Proof.
@@
 Lemma 2.9.
-Inequivalent $x$-colorings of $\Omega_{1}$ provide inequivalent $y$-colorings of $\Omega$.
+Fix $Y_{1},\ldots,Y_{x}$ colorings of $\Omega_{2}$, and let $X,X^{\prime}$ be colorings of $\Omega_{1}$.
+Let $Z$ and $Z^{\prime}$ be the colorings of $\Omega_{1}\times\Omega_{2}$ obtained from $X$ and $X^{\prime}$ respectively.
+If $X$ and $X^{\prime}$ are inequivalent, then $Z$ and $Z^{\prime}$ are inequivalent.
 Proof.
-Fix $Y_{1},\ldots,Y_{x}$ colorings of $\Omega_{2}$, and let $X,X^{\prime}$ be two colorings of $\Omega_{1}$.
-Let $Z$ and $Z^{\prime}$ be the colorings of $\Omega=\Omega_{1}\times\Omega_{2}$ obtained from $X$ and $X^{\prime}$ respectively.
 If $Z$ and $Z^{\prime}$ lies in the same orbit of $B\wr_{|\Omega_{1}|}A$, then the induced colorings of $Z$ and $Z^{\prime}$ on $\Omega_{1}$ need to be in the same orbit of $A$.
@@
 We work by induction on the cardinality of $\Omega$.
-Suppose that $\Omega_{1},\ldots,\Omega_{t}$ are the orbits of $G$ on $\Omega,$ and that we can find subsets $\Delta_{j}\subseteq\Omega_{j}$ such that $Stab_{G}(\Delta_{j})$ has orbits of length at most $6$ on $\Omega_{j}$, for each $j=1,\ldots,t$.
-Then $Stab_{G}(\sqcup_{j}\>\Delta_{j})$ will surely work.
+Suppose that $\Omega_{1},\ldots,\Omega_{t}$ are the orbits of $G$ on $\Omega,$ and that we can find subsets $\Delta_{j}\subseteq\Omega_{j}$ such that $\mathop{\mathrm{Stab}}_{G}(\Delta_{j})$ has orbits of length at most $6$ on $\Omega_{j}$, for each $j=1,\ldots,t$.
+Then $\mathop{\mathrm{Stab}}_{G}(\sqcup_{j}\>\Delta_{j})$ will surely work.
 It follows that we can suppose that $G$ is transitive.
-Primitive groups can be handled with Lemma 2.2 and Table 1.
+The base cases for the induction, primitive groups, can be handled with Lemma 2.2 and Table 1.
 Let $\Omega=\Omega_{1}\times\Omega_{2}$, with $|\Omega_{1}|\geq 2$.
@@
     Here $\ell_{2,2}(B)=3$.
-Using the $2$-asymmetric orbits of $B$, and the $3$-asymmetric $3$-colorings of $\Omega_{1}$, arguing as before we can construct $\ell_{3,3}(A)$ good orbits of $G$ (set $x=3$, $y=2$ in the construction above).
+Using the $2$-asymmetric orbits of $B$, and the $3$-asymmetric 33-colorings of $\Omega_{1}$, arguing as before we can construct $\ell_{3,3}(A)$ good orbits of $G$ (set $x=3$, $y=2$ in the construction above).
 In fact, note that orbit lengths are multiplied in the imprimitive action, so the length of each orbit will be bounded by $2\cdot 3=6$.
@@
     We can assume $B\cong\mathop{\mathrm{Sym}}(3)$, so that $\ell_{2,3}(B)=4$.
-Using the $3$-asymmetric orbits of $B$, and the $2$-asymmetric $4$-colorings of $\Omega_{1}$, as before we can construct $\ell_{4,2}(A)$ good orbits of $G$ (set $x=4$, $y=2$ in the construction above).
+Using the $3$-asymmetric orbits of $B$, and the $2$-asymmetric 44-colorings of $\Omega_{1}$, as before we can construct $\ell_{4,2}(A)$ good orbits of $G$ (set $x=4$, $y=2$ in the construction above).
 The length of each orbit will be bounded by $3\cdot 2=6$.
@@
 The cardinality of the subset in Theorem 1.1 grows linearly with the cardinality of $\Omega$, which is necessary as shown by $B\wr K$, when $B$ has some orbit of length exceeding $6$, and the degree of $K$ grows.
-The bound on the derived length follows easily:
 Proof of Corollary 1.2.
 Let $\Delta\subseteq\Omega$ be as provided by Theorem 1.1.
-Then $Stab_{G}(\Delta)$ can be embedded in a direct product of copies of $\mathop{\mathrm{Sym}}(6)$.
+Then $\mathop{\mathrm{Stab}}_{G}(\Delta)$ can be embedded in a direct product of copies of $\mathop{\mathrm{Sym}}(6)$.
 Every solvable subgroup of $\mathop{\mathrm{Sym}}(6)$ has derived length at most $3$, and the proof follows.
 ∎
-In a previous version of this paper, Corollary 1.2 was obtained directly, without first bounding the length of the orbits.
-Indeed, the idea of proving a stronger statement by induction and gluing colorings is flexible enough to give structural results directly (see Theorem 2.16 below, for example).
-The next example provides a transitive solvable group where no subset gives a metabelian stabilizer:
+The next example provides a transitive solvable group where no subset gives a metabelian stabilizer.
 Example 2.11.
@@
 Since $\mathop{\mathrm{Sym}}(3)\wr_{2}\mathop{\mathrm{Sym}}(2)$ is not supersolvable, neither is $H\wr_{2}\mathop{\mathrm{Sym}}(2)$, and the same example shows that supersolvable set-stabilizers need not to exist.
-This is an improvement on [Glu24, Example 3.4], which deals with nilpotent set-stabilizers.
+This is an improvement on [Glu25, Ex.
+3.4], which deals with nilpotent set-stabilizers.
 The next similar construction involves $\mathrm{A\Gamma L}_{1}(8)$, and provides a transitive solvable group where no subset gives a stabilizer with all orbits of length at most $5$.
@@
 We argue as in Subsection 2.3, with $x=5$ and $y=3$.
-Using the $2$-asymmetric $3$-colorings of $\Omega_{2}$ as colors for $\Omega_{1}$, and the asymmetric $5$-colorings of $\Omega_{1}$, we can construct $\ell_{5,1}(A)$ inequivalent $2$-asymmetric $3$-colorings of $\Omega$.
+Using the $2$-asymmetric 33-colorings of $\Omega_{2}$ as colors for $\Omega_{1}$, and the asymmetric $5$-colorings of $\Omega_{1}$, we can construct $\ell_{5,1}(A)$ inequivalent $2$-asymmetric 33-colorings of $\Omega$.
 We have $\ell_{5,1}(A)\geq 5$ from Corollary 2.6, which makes the induction smooth and completes the proof.
 ∎
-The correspondent structural result answers [Bab22, Question 7c].
+The correspondent structural result answers [Bab22, Quest. 7c].
 Corollary 2.14.
 Let $G$ be a solvable permutation group on the finite set $\Omega$. Then there exists a $3$-coloring of $\Omega$ whose stabilizer is an elementary abelian $2$-group.
-See [Glu24] for another positive result (with two colors) that emphatizes the importance of the prime $2$.
-In general, five colors are necessary and sufficient to achieve a trivial stabilizer [Ser96, Theorem 1.2].
+See [Glu25] for another positive result (with two colors) that emphasizes the importance of the prime 22.
+In general, five colors are necessary and sufficient to achieve a trivial stabilizer [Ser96, Th.
+1.2].
 2.5. CFSG-free results for nonsolvable groups
 The foundation stone to prove a result like Theorem 1.1 is the fact that a primitive solvable group of large degree has a regular orbit on the power set [Glu83].
-With the trivial exceptions of the alternating and symmetric groups, the same is actually true in the nonsolvable case, but the known proofs of Cameron, Neumann, Saxl and Seress [CNS84, Ser97] require the Classification of Finite Simple Groups.
+With the trivial exceptions of the alternating and symmetric groups, the same is true in the nonsolvable case, but the known proofs of Cameron-Neumann-Saxl and Seress [CNS84, Ser97] require the Classification of Finite Simple Groups.
 Actually, Cameron [Cam86] proved a stronger statement concerning the number of regular orbits.
 We observe that a more recent elementary result of Sun and Wilmes [SW15, SW16] is strong enough to provide an easier proof of Cameron’s theorem.
-[^1] Babai pointed out that a slightly weaker observation, which gives an elementary proof of [CNS84], was made by himself in [Bab22, Theorem 15.1].
+[^1] Babai pointed out that a slightly weaker observation, which gives an elementary proof of [CNS84], was made by himself in [Bab22, Th.
+15.1].
 Theorem 2.15.
-Let $G$ be a primitive group of degree $n$ not containing $\mathop{\mathrm{Alt}}(n)$. Then the probability that a random orbit on the power set is a regular orbit tends to $1$ as $n$ grows.
+Let $G$ be a primitive group of degree $n$ not containing $\mathop{\mathrm{Alt}}(n)$. Then the proportion of regular orbits on the power set tends to $1$ as $n\to\infty$.
 Proof.
-With the notation in [Cam86], it is sufficient to prove that $\varepsilon_{1}(G)\to 0$ when $n$ grows.
-We mimic Cameron’s proof, which uses the notions of base size and minimal degree in a clever way.
-At the crucial point, he obtains the desired result provided that $|G|\leq c^{\sqrt{n}}$, where $c$ is a constant (see the Remark at the end of page 308).
-Then, the Classification of Finite Simple groups is used to bound the order of a primitive group, remaining only with Johnson and Hamming groups, which are handled separately.
-While Babai’s theorem [Bab81] is not enough, [SW16, Corollary 1.6] is sufficient for this job, having a thresold of $d^{\,n^{1/3}(\log n)^{7/3}}$ where $d$ is a constant.
-∎
-We stress that Babai’s theorem [Bab81] has several applications, and Theorem 2.15 is a nice example where Sun-Wilmes improvement turns out to be crucial.
-(See [Ebe24] for another recent example.) We are now ready to prove Theorem 1.3, which answers [Bab22, Question 6a].
-Again, we prove a slightly stronger statement.
+By the remark near the bottom of page 308 of [Cam86], it suffices to show that $|G|\leq c^{\sqrt{n}}$ for some absolute constant $c$ - this is immediate from [SW16, Cor. 1.6]. ∎
+We stress that Babai’s theorem [Bab81] has several applications, and Theorem 2.15 is a nice example where Sun-Wilmes improvement turns out to be crucial. (See [Ebe24] for another recent example.)
+We are ready to prove Theorem 1.3. Again, we prove a slightly stronger statement.
 Theorem 2.16.
@@
 After Theorem 2.15, it is easy to see that such a constant $C$ exists for all primitive groups.
-Given $C$, we call good the orbits on the power set with the property in the statement.
+Given $C$, we call nice the orbits on the power set with the property in the statement.
 We can suppose that $G$ is transitive.
@@
 Let $G\leqslant B\wr_{|\Omega_{1}|}A$, with $A$ primitive on $\Omega_{1}$, and $B$ transitive on $\Omega_{2}$.
-By induction, $B$ admits two good orbits on $\mathcal{P}(\Omega_{2})$ (equivalently, two good $2$-colorings).
-Arguing as in Subsection 2.3 (set $x=y=2$), using these good orbits of $B$ as colors for $\Omega_{1}$, and the two good $2$-colorings of $\Omega_{1}$, we can construct two inequivalent good orbits of $G$ on $\mathcal{P}(\Omega)$.
+By induction, $B$ admits two nice orbits on $\mathcal{P}(\Omega_{2})$ (equivalently, two nice $2$-colorings).
+Arguing as in Subsection 2.3 (set $x=y=2$), using these nice orbits of $B$ as colors for $\Omega_{1}$, and the two nice $2$-colorings of $\Omega_{1}$, we can construct two inequivalent nice orbits of $G$ on $\mathcal{P}(\Omega)$.
 The proof is complete.
 ∎
-With more work, and using the Classification of Finite Simple Groups, it should be possible to find an explicit version of Theorem 2.16 (it is enough to find $C$ for all primitive groups).
+Using the Classification of Finite Simple Groups, it should be possible to find an explicit version of Theorem 2.16 (it is enough to find $C$ for all primitive groups).
 3. Primitive groups with solvable stabilizer
@@
 Therefore, we will focus on primitive groups, with the goal of proving Theorem 1.5.
-This proof is more algebraic, but the combinatorial ideas in the previous section will also play a key role.
-In the case of primitive solvable groups, Theorem 1.5 is [LS24, Theorem 1.1].
-This part is really about finding a good centralizer in every irreducible solvable linear group, which was done by Yang [Yan09] using ideas in [MW04].
-In general, as explained by Li and Zhang in [LZ11, Sec.
+This proof is more algebraic, but the combinatorial ideas in the previous section will play a key role.
+In the case of primitive solvable groups, Theorem 1.5 is [LS25, Th.
+1.1].
+This part is really about finding a good centralizer in irreducible solvable linear groups, which was done by Yang [Yan09] using ideas in [MW04].
+Using the O’Nan-Scott theorem, as explained by Li and Zhang in [LZ11, Sec.
 2], it is not hard to see that a primitive nonsolvable group with solvable stabilizer is almost simple or of a specific product type.
@@
 Remark 3.3.
-In an almost simple group $G$, the socle $T$ is the unique minimal normal subgroup.
-In fact, let $N\neq T$ be another minimal normal subgroup.
-Then $[N,T]=1$ and so $N\subseteq C_{G}(T)=1$.
-This implies that the solvable radical of $G$ is trivial and that every solvable subgroup of $G$ is core-free.
+In an almost simple group $G$, the socle $T$ is the unique minimal normal subgroup. This implies that the solvable radical of $G$ is trivial and that every solvable subgroup of $G$ is core-free.
 Proof of Theorem 1.4.
 Let $T$ be the socle of $G$, and let $G_{0}$ be the normal subgroup containing $T$ which is described in [LZ11, Pag.
-4].
-We have that the maximal subgroup $H$ is an extension of $H_{0}=H\cap G_{0}$ with
-$H/H_{0}\cong G_{0}H/G_{0}\cong G/G_{0},$
-which is a section of $G/T$ and so of $\mathop{\mathrm{Out}}(T)$. It follows that
+444].
+In particular, $G_{0}$ is minimal such that $H_{0}=H\cap G_{0}$ is a maximal subgroup of $G_{0}$.
+It follows that the maximal subgroup $H$ of $G$ is an extension of $H_{0}$ with
+$H/H_{0}\cong G_{0}H/G_{0}=G/G_{0},$
+which is a section of $G/T$ and so of $\mathop{\mathrm{Out}}(T)$. Therefore
 $dl(H)\leq dl(H_{0})+dl(G/G_{0}).$
@@
 This and Lemma 3.2 prove Theorem 1.4 for some absolute constant.
-The (relatively high) constant $10$ is attained by the Fischer group $Fi_{{}_{23}}$, which makes smooth the remaining computations in (i)-(iii).
-A more careful inspection of the tables shows that $dl(H_{0})\leq 6$ in the infinite families of (iii), while we use a computer for the remaining single groups.
-We point out some mistakes in [LZ11, Table 15], and rely instead on the [GAP4] library CTblLib (v1.3.9).
-[^2] One mistake is in correspondence of $Fi_{{}_{23}}$.
-In fact, $Fi_{{}_{23}}$ has a unique conjugacy class of solvable maximal subgroups, namely AtlasSubgroup(‘‘Fi23’’,7)), with structure $43^{1+8}.2^{1+6}.3^{1+2}.2S$ 4 (with the ATLAS notation) and derived length $10$.
+The (relatively high) constant $10$ is attained by the Fischer group $\mathop{\mathrm{Fi}}_{{}_{23}}$, which makes smooth the remaining computations in (i)-(iii).
+A more careful inspection of the tables shows that $dl(H_{0})\leq 6$ in the infinite families of (iii), while we use a computer for the remaining groups.
+We point out that there are some mistakes in [LZ11, Tab.
+15], and rely instead on the [GAP4] library CTblLib (v1.3.9).
+[^2] One mistake is in correspondence of $\mathop{\mathrm{Fi}}_{{}_{23}}$.
+In fact, $\mathop{\mathrm{Fi}}_{{}_{23}}$ has a unique conjugacy class of solvable maximal subgroups, namely AtlasSubgroup(‘‘Fi23’’,7)), with structure $3^{1+8}.2^{1+6}.3^{1+2}.2S_{4}$ (with the ATLAS notation) and derived length $10$.
 A potential obstruction is given by a certain conjugacy class of solvable maximal subgroups $H$ of the baby monster, which seems to be not available on a computer.
-With the ATLAS notation [CCN+85], $H$ is isomorphic to $3^{11}.(S4\times 2S4)$.
+With the ATLAS notation [CCN+85], $H$ is isomorphic to $3^{11}].(S_{4}\times 2S_{4})$.
 Any group of order $3^{11}$ has nilpotency class at most $10$ and derived length at most $4$, showing that $dl(H)\leq 8$ in this case.
 ∎
-A weaker version of Theorem 1.4 says that there exists a constant $C$ such that if $G$ is a finite group with a solvable maximal subgroup $H$, and the derived length of $H$ is more than $C$, then $G$ is not simple.
-It would be interesting to find a proof of this curious fact which does not rely on the Classification of Finite Simple Groups.
+A cheap version of Theorem 1.4 says that there exists a constant $C$ such that if $G$ is a finite group with a solvable maximal subgroup $H$, and the derived length of $H$ is more than $C$, then $G$ is not simple.
+It would be interesting to find an elementary proof of this curious fact.
 3.2. Groups of product type
@@
 Let $d\geq 2$, and let $S\leqslant\mathop{\mathrm{Sym}}(d)$ be a transitive solvable group.
-By [LZ11, Theorem 1.1], the primitive group $G$ that we care about satisfies
+By [LZ11, Th.
+1.1], the primitive group $G$ that we care about satisfies
 $T\wr_{d}S\>\leqslant\>G\>\leqslant\>X\wr_{d}S,$
@@
 Proof of Theorem 1.5.
+By the discussion above, it rests only to deal with the product type case.
 Let $V\cong X^{d}$ be the base subgroup of $X\wr_{d}S$, and let $H<G$ be the stabilizer of a point in the product action.
@@
 Let $\pi\colon G\to S$ be the natural projection, and let $v\in V$.
-Since $Ker(\pi)=G\cap V$, we have
+Since $\mathop{\mathrm{Ker}}(\pi)=G\cap V$, we have
 (3.1) $dl(I_{v})\>\leq\>dl(I_{v}\cap V)+dl(\pi(I_{v})).$
@@
 Now $dl(Y)\leq 10$ by Theorem 1.4, and the rest of the proof deals with $\pi(I_{v})$.
-It is not hard to see that $C_{S}(v)=S\cap S^{v}\subseteq\pi(I_{v})$, but in general $\pi(I_{v})$ can be quite bigger.
+It is not hard to see that $C_{S}(v)=S\cap S^{v}\subseteq\pi(I_{v})$, but in general $\pi(I_{v})$ can be much bigger.
 For every $w\in W$, $\sigma\in S$ and $v\in V$, an explicit computation gives
@@
 Let $c_{1},\ldots,c_{r}\in X$ be representatives of these double cosets ($r\geq 2$).
-For any $v\in V$, and $j\in\{1,\ldots,r\}$, let
-$\Delta_{j}=\{i\in\{1,\ldots,d\}:x_{i}\in Yc_{j}Y\}.$
+For any $v=(x_{1},\ldots,x_{d})\in V$, and $j\in\{1,\ldots,r\}$, let
+$\Delta_{j}=\Delta_{j}(v)=\{i\in\{1,\ldots,d\}:x_{i}\in Yc_{j}Y\}.$
 Of course the $\Delta_{j}$’s form a partition of $\{1,\ldots,d\}$. The equality in (3.2) shows that $\pi(I_{v})$ only depends on this partition, and in particular
-(3.3) $\pi(I_{v})\>=\>\cap_{j=1}^{r}Stab_{S}(\Delta_{j})\>=\>\cap_{j=1}^{r-1}Stab_{S}(\Delta_{j}).$
+(3.3) $\pi(I_{v})\>=\>\cap_{j=1}^{r}\mathop{\mathrm{Stab}_{S}}(\Delta_{j})\>=\>\cap_{j=1}^{r-1}\mathop{\mathrm{Stab}_{S}}(\Delta_{j}).$
 To construct $v$, we will use only two elements $c_{1}$ and $c_{2}$ (and so two sets $\Delta_{1}$ and $\Delta_{2}=\{1,\ldots,d\}\setminus\Delta_{1}$).
@@
 This is assured if we choose representatives $c_{i}$’s which are in $T$.
-We can surely choose $c_{1}=1$ and $c_{2}\in T\setminus\{1\}$, because otherwise $T\leqslant Y$, which is impossible because $Y$ is core-free.
+We can surely choose $c_{1}=1$ and $c_{2}\in T\setminus Y$, because otherwise $T\leqslant Y$, which is impossible because $Y$ is core-free.
 Finally, if $\Delta_{1}=\Delta\subseteq\{1,\ldots,d\}$ is as provided by Corollary 1.2, from (3.3) we get $dl(\pi(I_{v}))\leq 3$, concluding by (3.1) that $dl(I_{v})\leq 13$ as desired.
@@
 As explained, one reason is that we want $v\in G$, but there is also a different motivation.
-In fact, it is well known that the number $r$ of double cosets of $Y$ in $X$ equals the rank of the primitive group $X$ with solvable stabilizer $Y$, and there actually exist some doubly transitive groups of this type (as $\mathop{\mathrm{Sym}}(5)$, for example).
+In fact, it is well known that the number $r$ of double cosets of $Y$ in $X$ equals the rank of the primitive group $X$ with solvable stabilizer $Y$, and there actually exist various doubly transitive groups of this type.
 A more detailed analysis of almost simple groups, and the additional use of Corollary 2.14, might help to find the right constant in Theorem 1.5.
-In the case of primitive solvable groups, this lies somewhere between $4$ and $9$ (see [LS24, Remark 2.9]).
+In the case of primitive solvable groups, this lies somewhere between $4$ and $9$ [LS25, Rem.
+2.9].
 H.Y.
@@
 Let $G\leqslant\mathop{\mathrm{Sym}}(\Omega)$ be a primitive group with solvable stabilizer.
-Then there exists $\Delta\subseteq\Omega$, $|\Delta|=\delta$, such that $Stab_{G}(\Delta)$ is solvable of derived length bounded by an absolute constant.
+Then there exists $\Delta\subseteq\Omega$, $|\Delta|=\delta$, such that $\mathop{\mathrm{Stab}}_{G}(\Delta)$ is solvable of derived length bounded by an absolute constant.
 Proof.
@@
 If $\{x_{1},x_{2}\}\subseteq\Delta$, then $C_{G}(\Delta)$ is solvable of bounded derived length.
-Now $Stab_{G}(\Delta)/C_{G}(\Delta)\leqslant\mathop{\mathrm{Sym}}(4)$, and it follows that $Stab_{G}(\Delta)$ is solvable of bounded derived lenght.
+Now $\mathop{\mathrm{Stab}}_{G}(\Delta)/C_{G}(\Delta)\leqslant\mathop{\mathrm{Sym}}(4)$, and it follows that $\mathop{\mathrm{Stab}}_{G}(\Delta)$ is solvable of bounded derived length.
 ∎